### v1

# A Spatial-statistical model to analyse historical rutting data

Abstract
Pavement rutting poses a significant challenge in flexible pavements, necessitating costly asphalt resurfacing. To address this issue comprehensively, we propose an advanced Bayesian hierarchical framework of latent Gaussian models with spatial components. Our model provides a thorough diagnostic analysis, pinpointing areas exhibiting unexpectedly high rutting rates. Incorporating spatial and random components, and important explanatory variables like annual average daily traffic (traffic intensity), pavement type, rut depth and lane width, our proposed models account for and estimate the influence of these variables on rutting. This approach not only quantifies uncertainties and discerns locations at the highest risk of requiring maintenance, but also uncover spatial dependencies in rutting (millimetre/year). We apply our models to a data set spanning eleven years (2010-2020). Our findings emphasize the systematic unexpected spatial rutting effect, where more of the rutting variability is explained by including spatial components. Pavement type, in conjunction with traffic intensity, is also found to be the primary driver of rutting. Furthermore, the spatial dependencies uncovered reveal road sections experiencing more than 1 millimeter of rutting beyond annual expectations. This leads to a halving of the expected pavement lifespan in these areas. Our study offers valuable insights, presenting maps indicating expected rutting, and identifying locations with accelerated rutting rates, resulting in a reduction in pavement life expectancy of at least 10 years.
keywords:
Flexible pavement; predictive maintenance; rutting; spatial-statistical model; uncertainty
articletype: ARTICLE TEMPLATE
1 Introduction
Pavement maintenance is essential to ensure road safety and prevent excessive depreciation of roads. Rutting is a major problem in flexible pavement structures and it is one of the major causes why pavements need asphalt resurfacing. Rutting is a distress mechanism that causes longitudinal surface depressions within the wheel path of the asphalt concrete layer. It can substantially affect traffic safety (Hao et al. 2009) due to impaired steerability and friction reduction due to standing water in the ruts. Ruts occur under the combined actions of traffic loads (Zhang et al. 2017) and is affected by environmental conditions, e.g., climate, seepage, temperature, weather (Ranadive and Tapase 2016, Gudipudi et al. 2017, Zhang et al. 2023), and the inherent viscoelastic properties of the construction (e.g., thickness in surface layers or voids in the mixture Gu et al. 2017, Cheng et al. 2023). Rutting may occur at the various layers (Lea and Harvey 2015, Alhasan et al. 2018) or various stages (e.g., primary, secondary or tertiary) (Korkiala-Tanttu and Dawson 2007, Al-Khateeb and Basheer 2009, Gupta et al. 2014) of a constructed pavement. In cold climates, rutting often determines the longevity of the top layer of flexible pavements (Lundy et al. 1992), and the dominating mechanism is wear from studded tires (Snilsberg 2008), as opposed to permanent deformations in the mastic.
The rut depth can be effectively measured by laser profilometry (Thodesen et al. 2012). Many road agencies or road maintenance contractors map their road network regularly as input for their Pavement Management System (PMS). In Norway, all state roads are scanned every year, and this measurement program have been providing pavement condition data into a single database (ROSITA) since 2010. The Norwegian PMS system can provide automatically generated reports, for example which sections exceed the maximum allowed rut depth defined in the maintenance standards. However, deeper diagnostic analysis to identify areas that exhibit unexpectedly high rutting rates still requires manual labor. Such type of data analysis is typically only initiated based on suspicion and are very time consuming.
Acknowledging that the value of long time series of pavement condition data has not been fully utilized, the Norwegian Public Roads Administration initiated a research project that aims for more automated pavement diagnostics. Our working hypothesis is that advanced statistical analysis of the available data can provide a better understanding of how the pavement condition of a whole network develops and allow automated identification of segments that perform better or worse than expected. Moving from “manual, suspicion-based” diagnostics towards automated identification of problem areas enables pavement managers to concentrate their diagnostic resources on finding the correct countermeasures, rather than finding the problem areas. Identifying sections with unexpectedly high rutting levels requires that one compares with the ‘expected performance’ in terms of rutting.
Several empirical models have been proposed in the past to predict the performance of pavements (Hu et al. 2022, Yao et al. 2023, Yu et al. 2023) and these models typically include rutting. Empirical models usually relate the accumulation of distresses (e.g., rut, roughness, cracks) in the pavement layers to the load repetitions (annual average daily traffic) and other affecting factors (e.g., pavement or asphalt type, weather, width of the road, layer thickness). Empirical models are known to suffer from limited applicability (Gupta et al. 2014), so when the conditions (for example road construction, weather conditions or damage mechanisms) during the data collection differs from the conditions where the model is applied, one cannot expect to get reasonable results. Empirical models developed for pavement wear in cold climates exists (Jacobson and Wågberg 2005), but the challenge of limited predictability remains present. Calculating the expected rutting for sections based on an empirical model is therefore likely to rather large uncertainty.
Another modelling framework for predicting pavement performance from local historical data has been based on foundation layers and stages of deterioration and models of transition probability matrices (TPMs) (Chou et al. 2008, Pérez-Acebo et al. 2019, Yamany and Abraham 2020). This has some weaknesses, including its usability for the prediction of conditions based on a limited number of discrete classes (Vermunt et al. 1999, Pulugurta et al. 2009, Standfield et al. 2014) increasing imprecision and inaccuracy when assigning TPMs to the various deterioration stages; limitation in their generalizability to other flexible pavement data; and their lack of potential to account for spatial heterogeneity in the data (Saba et al. 2006, Chen and Mastin 2016)
Spatial models can provide additional explanation by accounting for trends in location. In the case of pavements, it is likely that sections close to each other behave more similar than sections that are far apart. These location components cannot explain the cause of the behaviour but they do have explanatory value and can be utilized to identify ‘hotspots’, predict future rut depths and quantify uncertainties.
Spatial models have been proposed for understanding pavement behaviour earlier (Lea and Harvey 2015, Alhasan et al. 2018, Ebrahimi et al. 2019). Lea and Harvey (2015) considered the applicability of spatial variability models on a test pavement containing a modified binder constructed for heavy vehicle simulator testing and found that the implementation of the spatial statistical techniques gave good insights concerning pavement variability following the complexity of pavement construction. (Alhasan et al. 2018), on the other hand, extended the more traditional mechanistic-empirical pavement prediction modelling framework by incorporating spatial variability and uncertainty in pavement foundation layers. (Ebrahimi et al. 2019) estimated asphalt surfacing lifetime for open road infrastructure on Norway road network (excluding bridges and tunnels) in Troms, Sør-Trøndelag and Vest-Agder. Their analysis showed that narrower road width was noted to result in shorter pavement lifetimes, particularly in Troms, where studded tires are used one month earlier than other locations, e.g., Vestland or Trøndelag in Norway.
Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced spatial rutting effects. To achieve this, we employ a hierarchical Bayesian framework comprising latent Gaussian models (LGMs) that encompass both random and spatial components. Latent Gaussian models have been used in various statistical analyses, as demonstrated by their applications in a multitude of fields (e.g., Dey et al. 2000, Chu et al. 2005, Banerjee et al. 2008, Rue et al. 2009b). These models offer a high degree of flexibility and explicitly account for model dependence. They can approximate entire processes defined over continuous domains and provide probabilistic predictions (Lindgren and Rue 2015). In our specific case, the rutting phenomenon on road surfaces persists regardless of whether it is directly observed, with the latent fields serving as the authentic representation of this rutting. A crucial premise underlying spatial models is that observations situated in close proximity to one another in the spatial domain exhibit greater similarity than those positioned further apart. For a broad class of LGMs, the Integrated Nested Laplace Approximation approach (INLA, Rue et al. 2009a) proves to be an apt tool for performing Bayesian inference.
The proposed models are based on historical rutting observations and a selection of explanatory variables known to influence pavement rutting. These variables include pavement type, Average Annual Daily Traffic (AADT), rut depth from the previous year and lane width. The inclusion of these variables is grounded in existing knowledge of factors that significantly influence rutting in road surfaces (Chan et al. 2010, Lang 2012, Thodesen et al. 2012, Ebrahimi et al. 2019, Yan et al. 2020, Zhao et al. 2020, Pan et al. 2021, Li et al. 2023). Rutting is a complex phenomenon, and the depth of rutting observed in the previous year is a key factor in predicting future rutting. Deeper rut depths from the previous year are expected to correlate with increased rutting in the current year, and this relationship is incorporated into the model. Additionally, lane width is considered as a contributing factor to rutting rates. It is logical to assume that narrower roads are more prone to higher rutting rates due to the limited space for motorists to maneuver. However, it’s important to note that the effect of lane width on rutting is contingent on the intensity of traffic. A critical element in our model is the role of Average Annual Daily Traffic (AADT) on the pavement. Rutting, as we argue, is contingent on the presence of AADT or the load imposed by traffic. Without traffic loading, certain forms of pavement deterioration, such as cracks and roughness, may occur, but rutting is improbable. This hypothesis aligns with prior research on the prediction of pavement variability and deterioration, as illustrated in the work of Hao et al. (2009) and Vedvik (2021) (Hao et al. 2009, Vedvik 2021). In this context, we emphasize a linear approximation for data processing, with the constraint that it passes through the origin. Furthermore, our model is designed with a standardized data approach, where the intercept is set to zero by design, ensuring consistency in the methodology.
As proof of concept, our proposed models are applied to eleven years (2010-2020) of pavement rutting data for one road stretch; the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border. We identify areas with unexplained or accelerated rutting and determine which explanatory variables are most important to rutting development. A detailed description of the data used and data cleaning process is given in Section 2. Model notation, formulation and evaluation criteria are given in Section 3. Results and discussion are given in 4. Section 5 explains how road authorities can use the model, based on the results.
2 Materials
2.1 Data source
The data material used in this research is obtained from the Norwegian Road Database (NVDB) and the Rosita database. The NVDB is primarily developed by the Norwegian Public Roads Administration (NPRA), the road authority for national roads, including managing, researching, planning, building, and operating. In 2022, NVDB consisted of 93 500 kilometres of public roads, including national roads (state-owned, 10,500 kilometres), county roads (44 000 kilometres) and municipal roads (39 000 kilometres) (Stephansen 2022). The Rosita database, which stores extensive and calculated data from measuring systems (see Section 2.3), was made available for use in this study.
2.2 Area of study
[…]
2.3 Data
The available explanatory variables for expected rut depth includes, traffic load in terms of annual average daily traffic (AADT), pavement or asphalt type (three categories), rut depth from the previous year and the lane width. A summary of the data used can be found in Table 1 and Figure 1 and Supplementary Material LABEL:supplementary:datasummary.
Table 1: Nomenclature and variables
Terms and Variables Description/Value Rut depth ($d_{i,t}$, millimetre) Measurement for year $t$ for all $20$-metre stretch ($i$); mean $=12.11$ millimetres and sd $=6.78$ millimetres Lane width (${w_{i}}$, metre) Half (or quarter) of width of 2-lane (or $2-2$ lane) used; mean $=6.02$ millimetres and sd $=0.69$ millimetres Pavement type Asphalt concrete (Ac), asphalt gravel concrete (Agc) and stone mastic asphalt (Sma) AADT (vehicles/year) $568-18500$ vehicles, with half (or quarter) of AADT for 2-lane (or $2-2$ lane) road used $r_{i,t}$ (millimetre/year) Change in rut depth between year $t$ and $t-1$ (for $t=1,...,T$).
2.3.1 Rut depth and rutting
The rut depth is measured annually using a laser scanner device ViaPPS (Pavement Profile Scanner) — a measurement system developed by ViaTech in cooperation with NPRA (Peraka and Biligiri 2020). The annual rut depth measurement is used to calculate rutting (mm/year), the indicator variable used for assessing pavement performance. We denote the rut depth variable at road segment $i=1,...,n$ in year $t$ as $d_{i,t}$, and $r_{i,t}$ as rutting (mm/year). Rutting for year $t$ at segment $i$ is defined as
$r_{i,t}=d_{i,t}-d_{i,t-1}.$ (1)
In the case of repavement, rut depth is smaller the proceeding year resulting in large negative values of $r_{i,t}$. Therefore, following Vedvik (2021), large negative rutting values are removed;
$r_{i,t}=\begin{cases}r_{i,t},&\text{if }r_{i,t}\geq-d_{i,t}/2\\ \text{NaN }&\text{otherwise, }\end{cases}$ (2)
i.e., $r_{i,t}$ is kept if it is at least half the rut depth, or assigned NaN (not a number) otherwise. Some missing rutting data may occur due to uncertainty arising from the actual rut depth measurements or the data collection process (Figure 1, graph (b) shows large variability in rutting).
2.3.2 Annual Average Daily Traffic (AADT)
The annual average number of daily traffic (AADT) f is available for each road segment for each year. AADT ranges from 568-18500 between the years $2010-2020$, with more traffic close to Stjørdal and less towards Sweden. NPRA states that high-traffic roads have an AADT of more than 5000 vehicles and recommends maintenance at rut depth values exceeding $20$ millimetres, while low-traffic roads with AADT lower than 5000 require pavement treatment when rut depth exceeds $25$ millimetres. Up to $2.5$ kilometres along the road, a quarter of the AADT is used in the analysis or half of the AADT otherwise. In all analyses, AADT in ten thousands are used.
2.3.3 Pavement type
In Norway, asphalt concrete is used on all types of roads and in all traffic classes. The asphalt mixtures used on the European route E14 are; asphalt concrete mixtures (Ac) and asphalt gravel concrete (Agc), and a stone mastic asphalt mixture (Sma). These include stone mastic asphalt Sma is the stronger and densely graded asphalt mass with unmodified binders and is mainly used on high-traffic roads; and asphalt concrete (Ac) and asphalt gravel concrete (Agc), which are mainly used on low-traffic roads of up to 5000 vehicles, courtyards, sidewalks and bicycle paths. These are softer, dynamic pavement mixtures, with less stringent requirements for the nature and grading of the stone material compared with that of stone mastic asphalt. Asphalt concrete, is however, more stringent that Asphalt gravel concrete; it is a dense-graded asphalt used as a protective layer over the road surface, or laid over a layer of asphalt gravel concrete (d’Angelo et al. 2008, vegvesen Vegdirektoratet 2009, Taddesse 2013, Zhang et al. 2022).
2.3.4 Lane width
The road width is quite variable and depends on the number of lanes. Since one lane is considered for the purposes of analysis, a quarter of the measured width up to 2.5 kilometres from Stjørdal or half of the measured width for the rest of the road stretch is used.
(a) AADT and Pavement type
(b) Rutting
(c) Lane width
Figure 1: AADT together with pavement type (a), rutting (for years 2019-2020, mm/year — missing in red) (b), lane width (metre) (c).
2.4 Exploratory analysis of rutting
On average, over all stretches and years, rutting is $1.67$ mm/year; see Figure  2 (a). The density plots,  2 graph (b), show rutting for all segments for each year. Some years having higher rutting levels and are more widely distributed than others. The lines in graph (c) gives the correlation of rutting values with respect to distance (mean across all years, and two randomly selected years; 2011 and 2013) (Bachmaier and Backes 2011). It is computed from the observed rutting and shows how the rutting values are correlated with respect to distance between the corresponding segments. The distance (metre) is shown on the x-axis, and the calculated correlation between rutting values is on the y-axis. The graph shows some spatial dependencies of pavement rutting along the road. These exist for short distances of up to ($\sim$) $300$ metres (3 kilometres).
Refer to caption (a) Observed rutting Refer to caption (b) Density Refer to caption (c) Autocorrelation
Figure 2: Graph (a): Observed rutting for each year (2011-2020), observed mean rutting $1.67$ mm/year (       ), and the observed mean rutting across segments ( ). Graph (b): Rutting density plots, graph. Graph (c) shows how rutting values are correlated with distance. Spatial dependencies in rutting exist up to 3 kilometres.
3 Latent Gaussian models for Rutting
A Bayesian latent Gaussian model is a three-stage hierarchical model consisting of a likelihood, a latent Gaussian field, and hyperparameters. In this paper, the true rutting observations $\boldsymbol{\mathrm{r}}=(r_{1},...,r_{n})$ at location $i=1,2...,n$ in year $t$ is modelled as a latent Gaussian field, assumed to have a Gaussian prior with zero mean and a small number of hyperparameters ($<15$). We specify six (6) models with different explanatory variables included and with and without spatial and random effects.
3.1 Latent field: model for true rutting
We denote the observed rutting for segment $i$ in year $t$ as $r_{i,t}$. Let $\eta_{t}(s_{i})$ denote the true rutting at location $s_{i}$. We model the observed rutting as Gaussian with expected value $\eta_{t}(s_{i})$ and variance $\sigma^{2}_{\varepsilon}$. Hence, $\sigma^{2}_{\varepsilon}$ is the uncertainty of the Gaussian observations. The unstructured error terms $\varepsilon_{i,t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\mathrm{Normal}(0,\sigma^{2}_{\varepsilon})$ are independent and identically distributed (i.i.d). Further, $r_{i,t}$ is conditionally independent, given the latent field $\eta_{t}(s_{i})$.
3.1.1 Likelihood: observed rutting given true rutting
We first model the observed rutting, given the true rutting. This is referred to as the likelihood model in the Bayesian hierarchical model setting. The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as
$r_{i,t}\ |\ \eta_{t}(s_{i})\sim\mathrm{N}\left(\eta_{t}(s_{i}).\sigma^{2}_{\varepsilon}\right).$ (3)
We only consider models that have an interaction with AADT, and include a random yearly effect. It is known that more traffic gives more rutting, so we include traffic (AADT) in the model as an interaction with each of the explanatory variables. This means that for pavement type, we assume that the rutting is linear for all pavement types (double traffic means double rutting), but that the different pavement types can have different rates of deterioration. From the explanatory analysis (Figure 2), we found that the rutting seems to be on different levels for different years, and this might be explained by different weather conditions. To account for this, we include a random yearly effect in the model.
3.1.2 Non-spatial rutting model: based on explanatory variables and year effect
The model without spatial effects (non-spatial) has the linear predictor
$\centering\begin{split}\eta_{t}(s)&=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\beta_{\mathrm{d-1}}z_{\mathrm{d-1}(s,t)}+\beta_{\mathrm{w}}z_{\mathrm{w}(s,t)}+\gamma_{t}.\end{split}\@add@centering$ (4)
The model assumes that the level of expected rutting (mm/year) is influenced by an interaction effect of the annual average daily traffic (AADT) and some explanatory variables, including the level of rut depth from the previous year ($z_{\mathrm{d-1}}$), the lane width ($z_{\mathrm{w}}$) and pavement type. Technically, $z_{Ac}(s,t)$ is
$\begin{cases}\text{AADT (in ten thousands) if pavement at ($s,t$)is Ac,}\\ 0\ \text{if not Ac;}\\ \end{cases}$
similarly, for pavement types Agc and Sma, representing asphalt gravel concrete and stone mastic asphalt, respectively. The effect of lane width, say, on rutting is dependent on the traffic intensity. The random yearly effect is assumed to be distributed as $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},...,\gamma_{T}){\sim}\mathrm{Normal}(\boldsymbol{0}_{T},\ \tau_{\gamma}^{-1})$.
3.1.3 Model including spatial effects
There are several variables not included in our model (Equation (4)) that are known to influence rutting. Examples are variables related to weather, bearing capacity and proportion of heavy traffic load. Most of these variables are spatially dependent, and segments close in space are likely to be more similar than segments further a part. This can be accounted for and utilised by including spatial random effects. The spatial random effect is a Gaussian process (GP) along the road and can be interpreted as the part of the residual (the part of the model not explained by the explanatory variables; lane width, rut depth from the previous year and pavement type) that is spatially dependent. Further, it is reasonable that some of the spatial random effect is consistent over years (e.g., bearing capacity, side slope height, surface and base curvature and proportion of heavy traffic) while others vary between years (e.g., heavy rain events or more heavy traffic due to construction work). Therefore, we include both an annually varying spatial process $\xi_{t}(\boldsymbol{s})$ for year $t=1,...,T$, and a common spatial process $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component (Ingebrigtsen et al. 2015). The spatial rutting component is the spatial contribution to the total rutting that is not explained by the other variables e.g., variables in the non-spatial model in Equation (4). This contribution can either be positive or negative, depending on the other variables and either over- or underestimate the observed rutting. The linear predictor for the model including spatial effects is described as
$\eta_{t}(s)$ $=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\beta_{\mathrm{d-1}}z_{\mathrm{d-1}(s,t)}+\beta_{\mathrm{w}}z_{\mathrm{w}(s,t)}+$ (5)
$\gamma_{t}+\xi_{t}(s)+\omega(s),$
where the remaining model parameters are defined as that in Equation 4.
The common spatial process $\omega(\boldsymbol{s})$ in Equations (5) is the time constant variable that can give rutting not explained by known variables (or unexplained rutting). It is a Gaussian random process $\left\{\omega(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{D}\subseteq\mathbb{R}^{d}\right\}$ such that for any $n\geq 1$ and for each set of spatial locations $\left(s_{1},...,s_{n}\right)$ satisfies
$\omega(\boldsymbol{s})=\left\{{\omega}(s_{1}),...,{\omega}(s_{n})\right\}=\left({\omega}_{1},...,{\omega}_{n}\right)\sim\mathrm{Normal}\left(\boldsymbol{\mu}^{\omega},\Sigma^{\omega}\right),$ (6)
where $\boldsymbol{\mu}^{\omega}=\left\{\mu^{\omega}(s_{1}),...,\mu^{\omega}(s_{n})\right\}$ is the mean vector and $\Sigma^{\omega}_{ij}=\mathrm{Cov}\left\{\omega(s_{i}),\omega(s_{j})\right\}=C^{\omega}\left\{\omega(s_{i}),\omega(s_{j})\right\}$ are the elements of the covariance matrix defined by the Matérn stationary isotropic covariance function. Simply, this covariance is expressed as $C^{\omega}\left(d;\sigma^{2}_{\omega},\rho_{\omega}\right)$, where $d=||s_{i}-s_{j}||$ is the distance between two road segments; $\sigma^{2}_{\omega(s)}$ is the marginal variance of the spatial rutting component and $\rho_{\omega}$ is the spatial range for the spatial rutting component. These are the parameters to be estimated.
Similarly for the annually varying spatial field $\xi_{t}(\boldsymbol{s})$, the parameters to be estimated are $\sigma^{2}_{\xi_{t}}$ and $\rho_{t}$. This can be done following the approach as in Equation 6. Detail description of the Matérn covariance structure is found in (e.g., Blangiardo and Cameletti 2015, Krainski et al. 2018) and the Supplementary Material LABEL:supplementary:modelformula_SUPPLEMENTARY.
3.1.4 Other models for comparison
We consider simpler non-spatial and spatial models for comparison. These models include some of the known explanatory variables and are presented in Table 2. The most complex spatial model (in Equation 5) we refer to as Model 1, and it’s non-spatial version, Equation (4), as Model 4. These models include three explanatory variables.
Table 2: Statistical models fitted
NAME MODEL DESCRIPTION
SPATIAL
Model 1 $\eta_{t}(s)=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\beta_{\mathrm{d-1}}z_{\mathrm{d-1}(s,t)}+\beta_{\mathrm{w}}z_{\mathrm{w}(s,t)}+\gamma_{t}+\xi_{t}(s)+\omega(s)$
Model 2 $\eta_{t}(s)=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\beta_{\mathrm{d-1}}z_{\mathrm{d-1}(s,t)}+\gamma_{t}+\xi_{t}(s)+\omega(s)$
Model 3 $\eta_{t}(s)=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\gamma_{t}+\xi_{t}(s)+\omega(s)$
NON-SPATIAL
Model 4 $\eta_{t}(s)=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\beta_{\mathrm{d-1}}z_{\mathrm{d-1}(s,t)}+\beta_{\mathrm{w}}z_{\mathrm{w}(s,t)}+\gamma_{t}$
Model 5 $\eta_{t}(s)=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\beta_{\mathrm{d-1}}z_{\mathrm{d-1}(s,t)}+\gamma_{t}$
Model 6 $\eta_{t}(s)=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\gamma_{t}$
3.2 Prior models
The priors on the explanatory variables $\boldsymbol{\beta}=(\beta_{\mathrm{Ac}},\beta_{\mathrm{Agc}},\beta_{\mathrm{Sma}},\beta_{\mathrm{d-1}},\beta_{\mathrm{w}})$; and the hyperparameters for the random yearly effect $\gamma$ are assumed as follows
[…]
Vague priors are used, and we let the data inform about the parameters to be estimated. The random yearly effect is evident in density plots in Figure 2 (c), which shows the varying yearly effect of rutting. For easier interpretations, the priors for the spatial components $\xi_{t}(s)$ and $\omega(s)$ are set through the spatial range $\rho$ and marginal standard deviation $\sigma_{\omega}(s)$ of the GRFs with Matérn covariance function in Equation (LABEL:eqn:randge_sigma_2). These are acquired jointly following the penalized complexity (PC) framework (Simpson et al. 2017, Fuglstad et al. 2019). The prior for the range is set through the probability $\mathrm{Pr}(\rho<250)=0.15$, indicating that there is a 15 percent probability that the spatial range is less than 250 metres. For the marginal standard deviation, the prior is set through the probability Pr$(\sigma_{\omega(s)}>2.5)=0.05$, indicating that there is a 5 percent probability that the marginal standard deviation for the rutting is over 2.5 meters. These priors are the same for all the spatial fields.
3.3 Inference and implementation
In Bayesian statistics, inference is done by computing posterior distributions for the parameters of interest based on prior models (Section 3.2), the likelihood model in Equation (3) and the data. Our goal is to estimate the marginal posteriors of the spatial rutting component, the regression coefficients, the hyperparameters and the expected rutting (marginal likelihood), given the known explanatory variables. We use INLA (integrated nested Laplace approximation) (Rue et al. 2009a, 2017, Martino and Riebler 2019), which is implemented in the R package R-INLA to approximate these marginal distributions.
3.4 Model evaluation
Three model fit measures are available; the deviance Information Criterion (DIC) (Spiegelhalter et al. 2002), the Watanabe-Akaike Information criterion (WAIC) (Watanabe and Opper 2010, Gelman et al. 2014) and the log-likelihood (Hubin and Storvik 2016) can be obtained for each of them in order to select the best one. The DIC and WAIC are based on hierarchical modeling generalizations of the Akaike Information Criterion (Akaike 1974), and are widely used in Bayesian statistics to perform model comparison. The values are determined by the model fit and its complexity, and by a penalty applied for overfitting. Models with a high level of complexity and a high goodness-of-fit, and models with the lowest DIC and WAIC should be chosen.
4 Results and discussion
We fit the models in Section 3.1.4 to the data in Section 2.3. For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation.
The results from the spatial Model 1 and the non-spatial version Model 4 are used to demonstrate the models’ performance. Based on Model evaluation criteria (Table 5), the spatial and non-spatial models with more explanatory variables (Model 1, 3 or 4) are selected as the best model, with the highest goodness-of-fit (lowest DIC and WAIC). While Models 3 or 5 capture the random (and/or spatial) variability of pavement rutting, they do not necessarily explain more of the uncertainty. Therefore, the spatial and non-spatial Models 1 and 4 are considered the most appropriate model to infer pavement performance from rutting data and are used for further analysis. The results for the remaining models are given in Appendix A.
4.1 Estimated spatial rutting component
Using Model 1 and the estimates obtained from fitting this model, we can identify areas wit h high rutting rates that are not explained by known factors such as pavement type or lane width. This is given by the posterior mean of the spatial rutting component $\omega(s)$. The spatial rutting component $\hat{\omega}(s)$ is plotted in Figure 3.
We find that the rutting inferred from $\hat{\omega}(s)$ is quite variable along the road; graph (a). The mean unexplained rutting contribution from $\hat{\omega}(s)$ is $0.01$ mm/year and the maximum unexplained portion is $3.21$ mm/year. This means that some locations would have $3.21$ mm/year extra rutting compared with the mean. There are also several locations where rutting exceeds $2$ mm/year (the observed mean rutting is $1.67$ mm/year; see Figure 2, graph (a)). This suggests that there are unexplained or unaccounted influential factors of rutting in a given year, one of which could be the weather. Rutting of at least $2$ millimetres at some locations (upper panel), or in a given year (lower panel), is concerning, given the expected lifetime of the pavement and the recommended levels for treatment. If maintenance is to be carried out at $20-25$ millimetres rut depth levels, depending on traffic, and pavement designs are expected to last between $20-30$ years, then some parts of the road would survive at most $12.5$ years ($10$ years in high-traffic areas). The minimum unexplained portion of rutting from $\hat{\omega}(s)$ is $-1.72$ mm/year. This suggests that those locations would have $1.72$ mm/year less rutting compared with mean ($0.01$ mm/year). That is, better real-life rutting performance than expected by the model. The rutting from the spatial component do appear to be relatively systematic, with a difference between the highest and lowest rutting values of $4.93$ millimetres.
Including spatial components in the model allow us to explain more of the rutting variability and identify areas with high rutting rates not explained by known variables. This is confirmed in Figure 3 (b), where the rutting explained by the random yearly effect from the spatial model is at least $0.5$ mm/year less compared with the non-spatial model. We further investigate locations with unexpectedly high rutting rates in Section 4.2.
Figure 3: Upper panel: Spatial rutting component (or unexplained portion of rutting $\{\hat{{\omega}}(s)\}$), with approximate 95% credible interval. Lower panel: Rutting from the random yearly effect ($\hat{{\gamma}_{t}}$) from Model 1 ( ) and Model 4 (       ), with approximate 95 % credible interval.
4.2 Estimated rutting from the explanatory variables, random yearly effect and spatial rutting component
We can also use the spatial model (Model 1) to identify locations, termed ‘hot spots’, that are in the greatest need of maintenance (Figure 4). For illustration, the model is used to infer rutting from the coefficients of the explanatory variables $\hat{\beta}$; the mean of the posterior means of the random yearly effects $\hat{\gamma}$; and the posterior mean of the spatial rutting component $\hat{\omega}(s)$. This is both the explained and unexplained rutting. The random yearly effect $\gamma_{t}$ varies from year to year, so we use an estimate of the mean, i.e., $\hat{\gamma}$. In addition, $\hat{\sigma}_{\xi_{t}}$ varies annually, but it is quite small (see Supplementary Materials LABEL:supplementary:addional_results) so it is not considered in our illustration.
First, we estimate the rutting from the sum of the coefficients of the explanatory variables and the mean of the posterior means of the random yearly effects $\hat{\gamma}$ only. The results are from Model 1, so we denote this by: $\hat{\eta^{*}}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}$. Next, we estimate the rutting from the sum of the coefficients of the explanatory variables, the mean of the random yearly effect $\hat{\gamma}$, and posterior mean of the spatial rutting component $\hat{\omega}(s)$; we denote by $\hat{\eta}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}+\hat{\omega}(s)$. Since the different asphalt types; stone mastic asphalt (Sma), asphalt concrete (Ac) and asphalt gravel concrete (Agc), are laid at specific locations along E14, these are highlighted in various colours along the plot line $\hat{\eta^{*}}$. These results are shown in Figure 4.
The effect of pavement type on rutting (i.e., the line showing the three asphalt types; Sma, Ac and Agc) is much less variable (ranging from $0.96$ to $3.89$ mm/year) compared with the sum of the explanatory variables, the mean of random yearly effect and the spatial rutting component (ranging from $0.03$ to $4.68$ mm/year); see graph (a). This suggests that the explanatory variables and the random yearly effect alone do not capture or realistically represent the variability in the expected rutting. Accounting for both spatial and random yearly effects allow more of the rutting variability to be explained. These can be seen in Table 3 and Table 4 in Appendix A, where $\hat{\sigma}_{\omega}(s)$ explains $0.47$ and $\hat{\gamma}$ explains $0.72$ compared with the remaining uncertainty ($\hat{\sigma}_{\varepsilon}=$ $1.38$) from the spatial model and the non-spatial version ($\hat{\sigma}_{\varepsilon}=$ 1.54). Additionally, the expected lifetime at some sections of the pavement is at most $6$ years considering only the explanatory variables and the random yearly effect or $5$ years considering the spatial rutting component $\hat{\omega}(s)$. This expected lifetime is depended upon traffic (i.e., $20$ or $25$ millimetres recommended maintenance for high- or low-traffic areas). We also see that the expected rutting $\hat{\eta}$ can be high, where the rutting is explained by the explanatory variables; or it could be moderate, but the spatial rutting component $\hat{\omega}(s)$ is large. Compared with the observed mean rutting across segments, which ranges from $-0.05$ to $8.33$ mm/year (see Figure 2, graph (a)), the maximum expected rutting estimated from the model (i.e., $4.68$ mm/year) is approximately halved the maximum of the observed mean rutting.
In graph (b), the expected rutting from the sum of the explanatory variables, the random yearly effect and the spatial rutting component; lane width; and AADT (per ten thousand) are plotted for a short section of the road. This part of the road, from segments $1500-2700$ (or $30000-54000$ metres towards Sweden), is dominated by the weaker pavement type Agc and, as shown in graph (a), has spikes of accelerated rutting. When plotted against the AADT or lane width, we see that rutting is higher with increasing traffic and narrow road lanes. This large variability may, presumably, be a result of: (i) a change in asphalt type from asphalt concrete (Ac) to asphalt gravel concrete (Agc, see same road stretch in the graph (a)); (ii) the type of traffic in that area, e.g., heavy-traffic (vehicles $>3500$ kilograms); (iii) poor road design; or (iv) lack of timely and efficient pavement rehabilitation.
The observed mean rutting (Figure 2 (a) in Section 2.4) is estimated to be $1.67$ mm/year. From graph (a), we found that the expected rutting, $\hat{\eta}$, for 351 road segments exceed this mean, and 22 of these road segments gave estimated rutting that exceeds $3$ mm/year (graphs (c) and (d), respectively). These 22 locations suggest potential hot spots for accelerated rutting and would require urgent maintenance as the expected lifetime would be at most 8 years, depending on traffic.
(a) Expected rutting from Model 1. Dashed line is mean rutting $1.67$ mm/year.
(b) Expected rutting, AADT and lane width
Refer to caption (c) $\hat{\eta}>2$ mm Refer to caption (d) $\hat{\eta}>3$ mm.
Figure 4: Graph (a): Expected rutting from: i) the spatial rutting component $\hat{\omega}(s)$, the random yearly effects plus coefficients of all explanatory variables, interacting with AADT for Model 1 $\left\{\hat{\eta}\right\}$ and ii) from the sum of the explanatory variables and the of random yearly effect only $\left\{\hat{\eta^{*}}\right\}$, with each asphalt type highlighted. The effects are inferred for the year 2020. Graph (b): Expected rutting plotted against ADDT and lane width for a small stretch of E14 (closer to Sweden). Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year or $3$ mm/year, respectively.
4.2.1 Investigating areas with high rutting rates
The large spike in Figure (4) (a), showing high rutting rates, is further investigated. We look at the road segments 2100 - 2500 metres surrounding this spike. This area is the municipality of Meråker, just 20 kilometres west of Storlien. This stretch, as shown in Figure 4 (a), gave high expected rutting, reaching $4.68$ millimetres, with an expected lifetime of just $5$ years (low-traffic area with required maintenance at 25 millimetres rut depth). This accelerated rutting could be because of clogged drainage, flooding, missing side ditches (Figure 5), heavy-traffic or road designs with railguards installed close to the lane edge, allowing that traffic drives more within the same wheel tracks. Meråker has industry and agriculture, therefore, roadways are more subjected to heavy vehicles, including farm vehicles such as tractors (industry, wheel drive, row crop).
(a) Short section between EV14 S6D1 m6148-m8212 F1 with high rutting.
(b) Location EV14 S6D1 m8017 F1 in year 2020
(c) Location EV14 S6D1 m7971 F1 in year 2021
(d) Location EV14 S6D1 m7961 F1 in year 2022
Figure 5: Graph (a): Location with high rutting rates (road segments $2100-2500$ towards Sweden). Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or go to the next url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@337286,7036109,14/vegsystemreferanse:337312.447:7036064.784
To further investigate this road section with high rutting values, we consider the effect of the explanatory variables, random yearly effect, and the spatial rutting component. We explore the spatial Model 1 using various values of the explanatory variables to approximate the rutting. First, we use the true values in the data, i.e., the asphalt type, the lane width and rut depth measured at that location, denoted by $\hat{\eta}_{\mathrm{True}}$. Second, the asphalt type is changed from its true state of asphalt gravel concrete (Agc) to asphalt concrete (Ac) or stone mastic asphalt (Sma); represented by $\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$, respectively. Third, the ‘best’ (or ‘worse’) values are used, i.e., if pavement type with the strongest binder, Sma, or a rut depth value from the previous year of $0.01$ millimetres ($\sim 0$), or lane width of $6$ metres were used. This is denoted by $\hat{\eta}_{\mathrm{Best}}$. These values are chosen following the reasoning: ($\sim$) zero rut depth means no rutting; wider roads mean a slower rate of rutting. The values considered to be ‘worse’, resulting in accelerated rutting, include a rut depth from the previous year of 25 millimetres, lane width of 2.75 metres and asphalt gravel concrete (Agc) pavement type. This is denoted by $\hat{\eta}_{\mathrm{Worse}}$. These values are found in the E14 data set. The cases are plotted in Figure 6 (a), which gives the expected rutting from the sum of the explanatory variables, th yearly random effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$.
Using the worse values in the analysis, the expected rutting is at least 2.3 ($2.25-7.94$) mm/year; Figure 6, graph (a) ( ). This means that with low AADT and the recommended maintenance rut depth level of 25 millimetres, pavement deterioration would accelerate and treatment or new pavement would be required within $3$ to $11$ years after initial construction. The best covariate values ( ) gave at most $4.26$ mm/year of rutting, with the least rutting value ($0.21$ mm/year). And, the earliest maintenance activity would be required from around 6 years after construction or maintenance. Further, the best values ( ) gave similar estimates of annual rutting to that of using the stronger asphalt binder type Sma (stone mastic asphalt,  ), except at road segments $2279-2388$, where the estimated rutting from only changing the asphalt type from asphalt gravel concrete to stone mastic is lower than that of using the best values. This suggests that asphalt type is the driving factor of expected rutting. Pavement types asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) gave similar estimates of expected rutting. This result is expected because they are made with similar binder concentration strengths.
(a) Expected rutting at road segments $2100-2500$ (metres) towards Sweden.
(b) Expected rutting from adding each explanatory variable added sequentially.
(c) Bearing capacity (tonn).
(d) SCI/BCI.
(e) SCI/BCI for the whole E14.
Figure 6: Graph (a): Expected rutting inferred from the spatial rutting component $\hat{\omega}(s)$ plus the ‘best’ (or ‘worse’) explanatory values at road segments $2100-2500$ towards Sweden. The estimate $\hat{\eta}_{\mathrm{Agc}}$ is for pavement type Agc plus the random yearly effect and the spatial rutting component. Graph (b): expected rutting inferred from $\hat{\omega}(s)$ plus the random yearly effect $\hat{\gamma}$ and each explanatory variable added sequentially in the model. Graph (c): bearing capacity for year 2021 (tonn), with 16 tonn or greater indicating a good parameter. Graphs (d) and (e) give the ratios of surface curvature index (SCI) to base curvature index (BCI) for road segments 2100-2500 (metres) or the whole E14, respectively, with values below $12$ indicating a good curvature index.
The above results suggest pavement type as the driving explanatory variable for expected rutting (Figure 6), so we determine its importance on rutting by changing the asphalt type from the true one (i.e., asphalt gravel concrete, Agc) to the stronger asphalt type Sma (stone mastic asphalt) and keeping all other variables (e.g., lane width) as given in the data set. The remaining explanatory variables are then added sequentially in the model with pavement type, i.e., stacking. Pavement type (Agc or Sma) is first added to the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$. We denote this by: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$, respectively. Next, rut depth from the previous year is added, and it is represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Similarly, the model with lane width $\{\hat{\beta}_{w}z_{\mathrm{w(s,t)}}\}$ added is represented by $\hat{\tilde{\eta}}_{\mathrm{w}}$. The results are shown in Figure 6, graph (b).
Results from using stone mastic asphalt (Sma) are represented by the solid lines, and from using asphalt gravel concrete (Agc), the dotted lines. The results confirm that pavement type is the driving factor of rutting. If the stronger binder type Sma were laid at this stretch instead of the true pavement type Agc, expected rutting would decrease, reaching at most $4.35$ mm/year (compared with 4.68 mm/year). This is approximately 0.33 mm/year less rutting annually, with an expected lifetime of 6 years, i.e., 1 extra year.
We also look at the available data for bearing capacity (Figure 6, graph (c)), surface curvature index (SCI) and base curvature index (BCI) for road segments 2100-2500. The data is available for year 2021, but it is a good indicator of the construction quality and the soil to sustain pressure over the past decade; Figure 6, graph (c). For bearing capacity, values greater than $16$ tonns (       ) suggest a good foundation design. The ratio of SCI to BCI ((Figure 6, graphs (d) and (e)) gives an indication of the location of the deflection (deformation) in the pavement layer (surface or base). The data suggests good bearing capacity, and that high rutting may be a result of other influential factors, e.g., side slope height (data not available), lack of side ditches to effect water drainage or deformation in the sub-base layer developed in thawing periods. We note that the bearing capacity data was sampled during the summer and corrected for temperature. During the summer, the measurements can be higher due to dry conditions, compared to those in thawing periods with high moisture content. Permanent deformation could have occurred during thawing periods, resulting in high rutting. There is also the possibility of varying construction material quality at various parts of the road, even though the pavement type (asphalt gravel concrete) is the same for the entire stretch (road segments $1500-3355$). Therefore, we investigate the consistency in the pavement construction by looking at deflections in the surface layer and the subsoil or reinforcement layer. The ratio of SCI to BCI is a good indication of the strengths or weaknesses in the construction, with $\mathrm{SCI/BCI}<12$ (       ) indicating a good parameter value. If this ratio is constant, then the construction is considered even, otherwise, there are variations indicating deformation in the various layers. The data, suggests an uneven construction of E14, with deeper deflections in the base layer.
We conclude that pavement type is the most important contributing factor to rutting. So, stronger binder types are recommended to ensure longer expected lifetimes.
4.3 Estimated model parameters
The estimated parameters and 95% credible intervals (CI) for the spatial Model 1 and non-spatial Model 4 are reported in Table 3. Comparing the coefficients for the explanatory variables between spatial models and their non-spatial version, we observe that the spatial models have considerably larger estimates for the regression coefficients of pavement type (see also Figure A.2 in Appendix A). Further, we see that the variance of the ‘noise’ term ($\hat{\sigma}_{\varepsilon}$) is larger for the models without the spatial rutting component. One explanation for this result is that the spatial rutting component account for unobservable factors, e.g., weather that vary in space, the effect of studded tires and other influencing factors such as traffic type that are not included in the model. Varying weather patterns across Norway and towards the Swedish border will affect the function of surface layers of pavement construction. In the winter time, increased temperatures result in a shorter time of frozen ground and snow on the surface and changes in the number of freeze-thaw cycles, thus, negatively affecting lower-grade binder types such as asphalt gravel concrete (Agc) and leading to increasing rutting. Asphalt gravel concrete is laid from () $(\sim)\ 30$ kilometres towards the Swedish border (Figure 1 (a)); an area that lie in the sub-artic climate zone with long severe winters and short warm to cool summers Peel et al. (2007), Ketzler et al. (2021), and having poorly developed networks (rural E14 highway west of Östersund) Jenelius (2010). This area is also more likely to be traversed by heavy-traffic (exceeding $3500$ kilograms or $3.5$ tonnes) Hovi et al. (2018), Pinchasik et al. (2020), negatively affecting pavement performance. More of the expected rutting is also explained by rut depth and lane width from the spatial model compared with the non-spatial model. This can be seen by the smaller values of the ‘noise‘ terms ($\hat{\sigma}_{\varepsilon}$) and uncertainty in the random year effect ($\hat{\sigma}_{\gamma}$) from the spatial model (see Table 3 and Table 4 in Appendix A for the simpler models).
Table 3: Estimated model parameters with $95\%$ credible intervals
Spatial Model Non-spatial Model
Model 1 Model 4
Mean/Median 95% CI Mean/Median 95% CI
$\hat{\beta}_{\mathrm{Ac}}$ 7.13 [6.62, 7.65] 3.25 [2.81, 3.68]
$\hat{\beta}_{\mathrm{Agc}}$ 7.34 [6.92, 7.76] 3.21 [2.75, 3.67]
$\hat{\beta}_{\mathrm{Sma}}$ 3.63 [3.37, 3.90] 2.00 [1.81, 2.19]
$\hat{\beta}_{\mathrm{d-1}}$ 0.74 [0.66, 0.82] 1.08 [0.98, 1.18]
$\hat{\beta}_{\mathrm{w}}$0.10 [-0.14, -0.07] 0.05 [-0.02, 0.11]
$\hat{\sigma}_{\omega(s)}$ 0.47 [0.44, 0.51] - -
$\hat{\sigma}_{\gamma}$ 0.72 [0.62, 0.88] 1.38 [0.96, 2.16]
$\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.54 [1.53, 1.56]
4.3.1 Effects of the explanatory variables
Interacting with the AADT, which is spatially varying, the effect of asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) on rutting may be explained by: heavy-traffic ($>3.5$ tonnes), whose proportion changes along the road and poor network design as these are constant in time. Also, various parts of the pavement have several layers due to non-removal during maintenance, leading to variability in pavement performance. The uncertainty from asphalt concrete and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) are also larger (wider 95% credible interval) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), but accounting for spatial dependencies allows more of the variance to be explained. Rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes more important when the spatial variability in pavement rutting is not considered (Model 4, Table 3). Further, instead of explaining the variability in expected rutting by pavement type, the non-spatial models explain this variability between years (Table 3 and Figure 3 (b)). The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is small and negative. This suggests that narrower lane widths would give higher expected rutting.
Based on these results, we conclude that changes in the coefficients are large when spatial dependencies of factors that are not modelled because of non-observable data (e.g., weather) are considered. And, the uncertainty explained is greater (smaller $\hat{\sigma}_{\varepsilon}$) for the spatial model.
5 How road authorities can make use of this model
The spatial-statistical model serves as a potent tool for road authorities, offering a data-driven approach to enhance maintenance strategies, identify critical areas for intervention and optimize road infrastructure durability. This approach facilitates a more ultimate diagnosis of rutting, empowering authorities to make informed decisions for sustainable road infrastructure.
Road authorities can use the spatial-statistical model for historical rutting data in several ways to improve maintenance strategies and prioritize interventions. These include; i) the identification of ‘hot spots’ — the spatial model helps pinpoint hot spots where rutting is accelerating, indicating areas in urgent need of maintenance. By analyzing coefficients of factors known to influence rutting (e.g., asphalt type), random effects, and spatial components, authorities can prioritize resources for targeted interventions in these critical locations; ii) comprehensive rutting diagnostic and understanding — the model provides a more comprehensive diagnostic and understanding of rutting. Explanatory variables and random yearly effects alone may not capture the full variability, but the inclusion of spatial components enhances the model’s explanatory power. This comprehensive understanding allows road authorities to make informed decisions about maintenance priorities; iii) extended expected lifetime — considering both spatial and random yearly effects extends the expected lifetime estimation of road sections. This information is valuable for long-term planning, enabling authorities to proactively address maintenance needs and allocate resources efficiently; iv) pavement type considerations — the model highlights the significant impact of pavement type on rutting rates. Road authorities can use this information to make informed decisions about pavement materials during construction or rehabilitation, optimizing durability and extending the overall lifetime of the road infrastructure. The recommendation of stronger asphalt binders, such as stone mastic asphalt (Sma), to enhance pavement durability, provides practical guidance for selecting materials that contribute to longer-lasting road surfaces; v) investigation of contributing factors — in-depth investigations into high rutting areas reveal potential contributing factors beyond pavement type, such as drainage issues, heavy traffic, and road design flaws. Road authorities can address these factors to prevent or mitigate rutting, ensuring a more holistic approach to road maintenance; vi) spatial model advantages — comparisons with non-spatial models highlight the advantages of including spatial components. The larger estimates for pavement type coefficients in spatial models emphasize the importance of spatial factors in explaining rutting variations. This information reinforces the importance of using spatial models for accurate and reliable inference and predictions; and finally vii) data-driven decision-making — the estimated model parameters and credible intervals provide a robust foundation for data-driven decision-making. Road authorities can rely on these parameters to make informed choices regarding maintenance priorities, resource allocation, and infrastructure improvements.
6 Conclusion
This research provides valuable insights into the factors influencing pavement rutting and the measurement of rut depth. The study confirms that the selection of asphalt type significantly affects rutting performance, with rut-resistant mixtures proving to be more effective in mitigating rutting. Lane width also plays a role in pavement rutting. Narrower roads tend to experience more concentrated traffic loads, which can lead to increased rutting. Wider roads, on the other hand, distribute the traffic load over a larger area, reducing the likelihood of rut formation. The data analysis indicated that narrower road stretches experienced higher rut depths. This outcome can be attributed to increased traffic volumes and higher wheel loads on roads, leading to more significant rutting. The findings, therefore, stress the importance of considering lane width when designing and maintaining pavements to minimize rutting distress. The recommendation is to incorporate a life-cycle cost-benefit perspective in the road design and planning process to minimize the potential for rutting issues over time. Further research is encouraged to explore additional factors influencing rutting; some of which include bearing capacity (Hopkins et al. 1998) as inadequate bearing strength of soil subgrades will not withstand construction traffic loading; ditch depth information that allows outflow of water from pavement structure; and lane width. This would likely increase the credibility of research results and the applicability of models to other road networks. In addition, it would be worthwhile to evaluate the long-term performance of rut-resistant asphalt mixtures in various traffic and environmental conditions.
Acknowledgement
The authors are grateful to Jørn Vatn and Jochen Köhler at the Norwegian University of Science and Technology for invaluable discussions on analysis and interpretation.
Disclosure Statement
The authors report no declarations of interest.
Funding
This work was supported by the Norwegian University of Science and Technology in collaboration with Statens Vegvesen Project [grant number 9056608 – SMARTere Vedlikehold (2019-2023)].
Appendix A Inference from model parameters
Table 4 compares estimates from the proposed spatial and non-spatial models. The results show that the spatial models account for uncertainty in factors that vary in space, e.g., weather and other explanatory variables such as bearing capacity and roadside ditches not included in the model. Including more explanatory variables reduces the ‘noise‘ term ($\hat{\sigma}_{\varepsilon}$) and random yearly effect ($\hat{\sigma}_{\gamma}$).
Table 4: Estimated model parameters with $95\%$ credible intervals
SPATIAL MODELS Model 1 Model 2 Model 3 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 7.13 [6.62, 7.65] 7.19 [6.66, 7.72] 6.98 [6.44, 7.53] $\hat{\beta}_{\mathrm{Agc}}$ 7.34 [6.92, 7.76] 7.38 [6.95, 7.81] 7.15 [6.72, 7.59] $\hat{\beta}_{\mathrm{Sma}}$ 3.63 [3.37, 3.90] 3.57 [3.30 3.85] 3.36 [3.08, 3.64] $\hat{\beta}_{\mathrm{d-1}}$ 0.74 [0.66, 0.82] 0.77 [0.69, 0.85] - - $\hat{\beta}_{\mathrm{w}}$0.10 [-0.14, -0.07] - - - $\hat{\sigma}_{\omega(s)}$ 0.47 [0.44, 0.51] 0.48 [0.45, 0.53] 0.52 [0.48, 0.57] $\hat{\sigma}_{\gamma}$ 0.72 [0.62, 0.88] 0.81 [0.64, 1.09] 0.72 [0.58, 0.94] $\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.38 [1.37, 1.39] 1.39 [1.38, 1.40] NON-SPATIAL MODELS Model 4 Model 5 Model 6 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 3.25 [2.81, 3.68] 3.25 [2.81, 3.69] 2.72 [2.28, 3.16] $\hat{\beta}_{\mathrm{Agc}}$ 3.21 [2.75, 3.67] 3.20 [2.74, 3.67] 2.71 [2.24, 3.18] $\hat{\beta}_{\mathrm{Sma}}$ 2.00 [1.81, 2.19] 2.02 [1.83, 2.21] 1.55 [1.36, 1.73] $\hat{\beta}_{\mathrm{d-1}}$ 1.08 [0.98, 1.18] 1.08 [0.98, 1.18] - - $\hat{\beta}_{\mathrm{w}}$ 0.05 [-0.02, 0.11] - - - - $\hat{\sigma}_{\gamma}$ 1.38 [0.96, 2.16] 1.45 [0.98, 2.21] 1.43 [0.99, 2.25] $\hat{\sigma}_{\varepsilon}$ 1.54 [1.53, 1.56] 1.54 [1.53, 1.56] 1.56 [1.54, 1.57]
A.1 Model Comparison
Model comparison criteria (Table 5) select the models with more explanatory variables most often as the best model, with the highest goodness-of-fit (lowest DIC and WAIC). As shown in the results above (e.g., Table 4), the more complex spatial models (Model 1) and non-spatial models (Model 4 or 5) capture the random and spatial variability of pavement rutting and explains more of the unstructured error terms with the addition of explanatory variable lane width.
Table 5: Model comparison
Model Estimates from spatial models Model Estimates from non-spatial models
DIC WAIC (-) loglikelihood DIC WAIC (-) loglikelihood
1 193678.05 287136.55 -101455.25 4 104290.62 100995.92 -52229.45
2 194145.55 301922.59 -101459.74 5 104292.58 101057.74 -52223.40
3 193938.88 285335.82 -101625.45 6 104763.18 101615.56 -52453.58
A.2 Marginal Posterior Distributions of Explanatory Variables
Figure A.2 gives the marginal posterior distributions of asphalt type, rut depth from the previous year and lane width. Narrow bands of the distributions indicate that the uncertainty is smaller. Graph (a) shows a clear shift in asphalt type’s contribution to rutting from the spatial and non-spatial models. This is because asphalt type is spatially varying. Stone mastic asphalt (Sma) gave lower rutting values because it has a stronger binder.
(a) Pavement type
(b) Rut depth
Figure A.2: Marginal posterior distributions of pavement type (solid lines are the spatial models, and the dotted lines are the non-spatial models), rut depth from the previous year and lane width.
A.3 Marginal Posterior Distributions of Random Effects
Figure A.3 gives the marginal posterior distributions of the standard deviation of the Gaussian observations ($\hat{\sigma}_{\varepsilon}$), the standard deviation of the random yearly effect ($\hat{\sigma}_{\gamma}$), the range of the common spatial field $\hat{\rho}$, and the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$. The solid lines represent the spatial models and the dotted lines represent the non-spatial models. Narrower densities suggest smaller estimated uncertainty. We find that some of the uncertainty of the Gaussian observations $(\hat{\sigma}_{\varepsilon})$ and random yearly effect $(\hat{\sigma}_{\gamma})$ go into the spatial effect $\{\hat{\sigma}_{\omega(s)}\}$, and this can be interpreted as some of the unobserved variability (e.g., weather, temperature) being explained by the spatial component; see Table 3 and Figure A.3 (a) and (b). This is confirmed by the smaller estimates of the ‘noise’ term $(\hat{\sigma}_{\varepsilon})$ compared with the non-spatial model (Table 4).
We examine the effects of the range of the common spatial field ($\hat{\rho}$), and the standard deviation of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$. As more explanatory variables are added to model (e.g., rut depth), the variability of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$ decreases; see Figure A.3, graph (d).
In Model 2, rut depth from the previous year is included, and this effect is large with the credible interval not covering zero. This indicates that deeper rut depths from the previous year would result in increased rutting; Figure A.2, graph (b). In addition, the spatial range for the common spatial field is, generally, larger for Model 2 compared with Model 3 (the simpler model), where rut depth is included (see Table LABEL:tab:range_for_each_year_appendix in Supplementary Material LABEL:supplementary:addional_results) for estimated $\hat{\rho}_{t}$ for each year). This is so, presumably, because of influenced factors such as varying weather patterns. But including lane width reduces the effect of rut depth on expected rutting.
Refer to caption (a) Marginal posterior distributions of the standard deviation of the Gaussian observations $\hat{\sigma}_{\varepsilon}$. Refer to caption (b) Marginal posterior distributions of the standard deviation of the random yearly effect $\hat{\sigma}_{\gamma}$. Refer to caption (c) Marginal posterior distributions of the nominal range of the common spatial field $\hat{\rho}$ Refer to caption (d) Marginal posterior distributions of the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$.
Figure A.3: Estimates of marginal posterior distributions of the spatial models (solid lines) and non-spatial models (dotted lines).

### v2

@@
 We apply our models to a data set spanning eleven years (2010-2020).
-Our findings emphasize the systematic unexpected spatial rutting effect, where more of the rutting variability is explained by including spatial components.
+Our findings emphasize the systematic unexplained spatial rutting effect, where some of the rutting variability is accounted for by spatial components.
 Pavement type, in conjunction with traffic intensity, is also found to be the primary driver of rutting.
@@
 Empirical models developed for pavement wear in cold climates exists (Jacobson and Wågberg 2005), but the challenge of limited predictability remains present.
-Calculating the expected rutting for sections based on an empirical model is therefore likely to rather large uncertainty.
+Calculating the expected rutting for sections based on an empirical model is therefore likely to give large uncertainty.
 Another modelling framework for predicting pavement performance from local historical data has been based on foundation layers and stages of deterioration and models of transition probability matrices (TPMs) (Chou et al.
@@
 2006, Chen and Mastin 2016)
-Spatial models can provide additional explanation by accounting for trends in location.
+Spatial models can provide additional explanation by accounting for trends in location, i.e., along the road.
 In the case of pavements, it is likely that sections close to each other behave more similar than sections that are far apart.
@@
 Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced spatial rutting effects.
-To achieve this, we employ a hierarchical Bayesian framework comprising latent Gaussian models (LGMs) that encompass both random and spatial components.
+To achieve this, we employ a hierarchical Bayesian framework comprising latent Gaussian models (LGMs) that encompass both random and spatial components (Jourdain and Klein-Paste 2024).
 Latent Gaussian models have been used in various statistical analyses, as demonstrated by their applications in a multitude of fields (e.g., Dey et al.
@@
 2009a) proves to be an apt tool for performing Bayesian inference.
-The proposed models are based on historical rutting observations and a selection of explanatory variables known to influence pavement rutting.
-These variables include pavement type, Average Annual Daily Traffic (AADT), rut depth from the previous year and lane width.
+The proposed models are based on historical rutting observations and a selection of explanatory variables known to influence pavement rutting that is available for the Norwegian stretch of the road.
+These variables are pavement type, Average Annual Daily Traffic (AADT), rut depth from the previous year and lane width.
 The inclusion of these variables is grounded in existing knowledge of factors that significantly influence rutting in road surfaces (Chan et al.
@@
 2023).
-Rutting is a complex phenomenon, and the depth of rutting observed in the previous year is a key factor in predicting future rutting.
+A critical element in our model is the role of Average Annual Daily Traffic (AADT) on the pavement.
+Rutting, as we argue, is contingent on the presence of AADT or the load imposed by traffic.
+Without traffic loading, certain forms of pavement deterioration, such as cracks and roughness, may occur, but rutting is improbable.
+This hypothesis aligns with prior research on the prediction of pavement variability and deterioration, as illustrated in the work of Hao et al.
+(2009) and Vedvik (2021).
+Rutting is a complex phenomenon, and the depth of rutting observed in the previous year is a key factor in predicting future rutting (Archilla and Madanat 2000).
 Deeper rut depths from the previous year are expected to correlate with increased rutting in the current year, and this relationship is incorporated into the model.
@@
 However, it’s important to note that the effect of lane width on rutting is contingent on the intensity of traffic.
-A critical element in our model is the role of Average Annual Daily Traffic (AADT) on the pavement.
-Rutting, as we argue, is contingent on the presence of AADT or the load imposed by traffic.
-Without traffic loading, certain forms of pavement deterioration, such as cracks and roughness, may occur, but rutting is improbable.
-This hypothesis aligns with prior research on the prediction of pavement variability and deterioration, as illustrated in the work of Hao et al.
-(2009) and Vedvik (2021) (Hao et al.
-2009, Vedvik 2021).
-In this context, we emphasize a linear approximation for data processing, with the constraint that it passes through the origin.
-Furthermore, our model is designed with a standardized data approach, where the intercept is set to zero by design, ensuring consistency in the methodology.
+Pavement type also plays a crucial role, particularly in colder climates where it can significantly influence rutting resistance.
+During winter, rising temperatures lead to shorter periods of frozen ground and snow cover on road surfaces, altering the frequency of freeze-thaw cycles.
+Consequently, this poses challenges for lower-grade binder types like asphalt gravel concrete (Agc), which are more susceptible to rutting under such conditions
 As proof of concept, our proposed models are applied to eleven years (2010-2020) of pavement rutting data for one road stretch; the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border.
@@
 Model notation, formulation and evaluation criteria are given in Section 3.
-Results and discussion are given in 4.
+Results and discussion are given in Section 4.
 Section 5 explains how road authorities can use the model, based on the results.
@@
 $r_{i,t}=d_{i,t}-d_{i,t-1}.$ (1)
-In the case of repavement, rut depth is smaller the proceeding year resulting in large negative values of $r_{i,t}$. Therefore, following Vedvik (2021), large negative rutting values are removed;
+In the case of repavement, rut depth is smaller in the proceeding year resulting in large negative values of $r_{i,t}$. Therefore, following Vedvik (2021), large negative rutting values are removed;
 $r_{i,t}=\begin{cases}r_{i,t},&\text{if }r_{i,t}\geq-d_{i,t}/2\\ \text{NaN }&\text{otherwise, }\end{cases}$ (2)
 i.e., $r_{i,t}$ is kept if it is at least half the rut depth, or assigned NaN (not a number) otherwise.
-Some missing rutting data may occur due to uncertainty arising from the actual rut depth measurements or the data collection process (Figure 1, graph (b) shows large variability in rutting).
+Some missing rutting data may occur due to uncertainty arising from the actual rut depth measurements or the data collection process (Figure 1, graph (c) shows large variability in rutting).
 2.3.2 Annual Average Daily Traffic (AADT)
@@
 Since one lane is considered for the purposes of analysis, a quarter of the measured width up to 2.5 kilometres from Stjørdal or half of the measured width for the rest of the road stretch is used.
-(a) AADT and Pavement type
+(a) AADT/Pavement type, and lane width
 (b) Rutting
-(c) Lane width
-Figure 1: AADT together with pavement type (a), rutting (for years 2019-2020, mm/year — missing in red) (b), lane width (metre) (c).
+Figure 1: AADT together with pavement type, and lane width (a), and rutting (for years 2019-2020, mm/year — missing in red) (b).
 2.4 Exploratory analysis of rutting
-On average, over all stretches and years, rutting is $1.67$ mm/year; see Figure  2 (a).
+On average, over all stretches and years, rutting is $1.62$ mm/year; see Figure  2 (a).
 The density plots,  2 graph (b), show rutting for all segments for each year.
 Some years having higher rutting levels and are more widely distributed than others.
-The lines in graph (c) gives the correlation of rutting values with respect to distance (mean across all years, and two randomly selected years; 2011 and 2013) (Bachmaier and Backes 2011).
+The lines in graphs (c) and (d) give the correlation of rutting values with respect to distance (mean across all years, and two randomly selected years; 2011 and 2012, respectively) (Bachmaier and Backes 2011).
 It is computed from the observed rutting and shows how the rutting values are correlated with respect to distance between the corresponding segments.
@@
 These exist for short distances of up to ($\sim$) $300$ metres (3 kilometres).
-Refer to caption (a) Observed rutting Refer to caption (b) Density Refer to caption (c) Autocorrelation
-Figure 2: Graph (a): Observed rutting for each year (2011-2020), observed mean rutting $1.67$ mm/year (       ), and the observed mean rutting across segments ( ).
+Refer to caption (a) Observed rutting Refer to caption (b) Density Refer to caption (c) Autocorrelation across years Refer to caption (d) Autocorrelation for years 2011-2012
+Figure 2: Graph (a): Observed rutting for each year (2011-2020), observed mean rutting $1.62$ mm/year (     ), and the observed mean rutting across segments ( ).
 Graph (b): Rutting density plots, graph.
-Graph (c) shows how rutting values are correlated with distance.
+Graphs (c) and (d) show how rutting values are correlated with distance.
 Spatial dependencies in rutting exist up to 3 kilometres.
@@
 A Bayesian latent Gaussian model is a three-stage hierarchical model consisting of a likelihood, a latent Gaussian field, and hyperparameters.
-In this paper, the true rutting observations $\boldsymbol{\mathrm{r}}=(r_{1},...,r_{n})$ at location $i=1,2...,n$ in year $t$ is modelled as a latent Gaussian field, assumed to have a Gaussian prior with zero mean and a small number of hyperparameters ($<15$).
+In this paper, the true rutting at location $s_{i}$ in year $t$ is denoted by $\eta_{t}(s_{i})$ and it is modelled as a latent Gaussian field that includes explanatory variables and spatial effects.
 We specify six (6) models with different explanatory variables included and with and without spatial and random effects.
+3.1 Likelihood: observed rutting given true rutting
+We denote the observed rutting for segment $i$ in year $t$ as $r_{i,t}$. The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as
+$r_{i,t}\ |\ \eta_{t}(s_{i})\sim\mathrm{N}\left(\eta_{t}(s_{i}),\sigma^{2}_{\varepsilon}\right).$ (3)
+Here, $\sigma^{2}_{\varepsilon}$ is the uncertainty of the Gaussian observations.
+The unstructured error terms $\varepsilon_{i,t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\mathrm{Normal}(0,\sigma^{2}_{\varepsilon})$ are independent and identically distributed (i.i.d).
 3.2 Latent field: model for true rutting
-We denote the observed rutting for segment $i$ in year $t$ as $r_{i,t}$.
-Let $\eta_{t}(s_{i})$ denote the true rutting at location $s_{i}$.
 We model the observed rutting as Gaussian with expected value $\eta_{t}(s_{i})$ and variance $\sigma^{2}_{\varepsilon}$.
-Hence, $\sigma^{2}_{\varepsilon}$ is the uncertainty of the Gaussian observations.
-The unstructured error terms $\varepsilon_{i,t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\mathrm{Normal}(0,\sigma^{2}_{\varepsilon})$ are independent and identically distributed (i.i.d).
 Further, $r_{i,t}$ is conditionally independent, given the latent field $\eta_{t}(s_{i})$.
-3.1.1 Likelihood: observed rutting given true rutting
-We first model the observed rutting, given the true rutting.
-This is referred to as the likelihood model in the Bayesian hierarchical model setting.
-The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as
-$r_{i,t}\ |\ \eta_{t}(s_{i})\sim\mathrm{N}\left(\eta_{t}(s_{i}).\sigma^{2}_{\varepsilon}\right).$ (3)
 We only consider models that have an interaction with AADT, and include a random yearly effect.
@@
 The random yearly effect is assumed to be distributed as $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},...,\gamma_{T}){\sim}\mathrm{Normal}(\boldsymbol{0}_{T},\ \tau_{\gamma}^{-1})$.
-3.1.3 Model including spatial effects
+3.2.2 Models including spatial effects
 There are several variables not included in our model (Equation (4)) that are known to influence rutting.
@@
 This can be accounted for and utilised by including spatial random effects.
-The spatial random effect is a Gaussian process (GP) along the road and can be interpreted as the part of the residual (the part of the model not explained by the explanatory variables; lane width, rut depth from the previous year and pavement type) that is spatially dependent.
+The spatial random effect is a Gaussian process (GP) along the road and can be interpreted as the part of the model not explained by the explanatory variables; lane width, rut depth from the previous year and pavement type that is spatially dependent.
 Further, it is reasonable that some of the spatial random effect is consistent over years (e.g., bearing capacity, side slope height, surface and base curvature and proportion of heavy traffic) while others vary between years (e.g., heavy rain events or more heavy traffic due to construction work).
-Therefore, we include both an annually varying spatial process $\xi_{t}(\boldsymbol{s})$ for year $t=1,...,T$, and a common spatial process $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component (Ingebrigtsen et al.
-2015).
+Therefore, we include both an annually varying spatial process $\xi_{t}(\boldsymbol{s})$ for year $t=1,...,T$, and a common spatial process $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component.
 The spatial rutting component is the spatial contribution to the total rutting that is not explained by the other variables e.g., variables in the non-spatial model in Equation (4).
-This contribution can either be positive or negative, depending on the other variables and either over- or underestimate the observed rutting.
+This contribution can either be positive or negative.
 The linear predictor for the model including spatial effects is described as
@@
 Similarly for the annually varying spatial field $\xi_{t}(\boldsymbol{s})$, the parameters to be estimated are $\sigma^{2}_{\xi_{t}}$ and $\rho_{t}$.
-This can be done following the approach as in Equation 6.
 Detail description of the Matérn covariance structure is found in (e.g., Blangiardo and Cameletti 2015, Krainski et al.
@@
 The most complex spatial model (in Equation 5) we refer to as Model 1, and it’s non-spatial version, Equation (4), as Model 4.
-These models include three explanatory variables.
+These models include all three explanatory variables.
 Table 2: Statistical models fitted
@@
 The prior for the range is set through the probability $\mathrm{Pr}(\rho<250)=0.15$, indicating that there is a 15 percent probability that the spatial range is less than 250 metres.
-For the marginal standard deviation, the prior is set through the probability Pr$(\sigma_{\omega(s)}>2.5)=0.05$, indicating that there is a 5 percent probability that the marginal standard deviation for the rutting is over 2.5 meters.
+For the marginal standard deviation, the prior is set through the probability Pr$(\sigma_{\omega(s)}>2.5)=0.05$, indicating that there is a 5 percent probability that the marginal standard deviation for the rutting is over 2.5 millimetres.
 These priors are the same for all the spatial fields.
@@
 The values are determined by the model fit and its complexity, and by a penalty applied for overfitting.
-Models with a high level of complexity and a high goodness-of-fit, and models with the lowest DIC and WAIC should be chosen.
+Models should balance complexity and goodness-of-fit, and models with the lowest DIC and WAIC should be chosen.
 4 Results and discussion
-We fit the models in Section 3.1.4 to the data in Section 2.3. For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation.
+We fit the models in Section 3.2.3 to the data in Section 2.3.
+For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation.
 The results from the spatial Model 1 and the non-spatial version Model 4 are used to demonstrate the models’ performance.
-Based on Model evaluation criteria (Table 5), the spatial and non-spatial models with more explanatory variables (Model 1, 3 or 4) are selected as the best model, with the highest goodness-of-fit (lowest DIC and WAIC).
-While Models 3 or 5 capture the random (and/or spatial) variability of pavement rutting, they do not necessarily explain more of the uncertainty.
-Therefore, the spatial and non-spatial Models 1 and 4 are considered the most appropriate model to infer pavement performance from rutting data and are used for further analysis.
+Based on Model evaluation criteria (Table 5), and for easy comparison of the models, the spatial and non-spatial models with more explanatory variables are selected.
 The results for the remaining models are given in Appendix A.
-4.1 Estimated spatial rutting component
-Using Model 1 and the estimates obtained from fitting this model, we can identify areas wit h high rutting rates that are not explained by known factors such as pavement type or lane width.
-This is given by the posterior mean of the spatial rutting component $\omega(s)$.
-The spatial rutting component $\hat{\omega}(s)$ is plotted in Figure 3.
-We find that the rutting inferred from $\hat{\omega}(s)$ is quite variable along the road; graph (a).
-The mean unexplained rutting contribution from $\hat{\omega}(s)$ is $0.01$ mm/year and the maximum unexplained portion is $3.21$ mm/year.
-This means that some locations would have $3.21$ mm/year extra rutting compared with the mean.
-There are also several locations where rutting exceeds $2$ mm/year (the observed mean rutting is $1.67$ mm/year; see Figure 2, graph (a)).
-This suggests that there are unexplained or unaccounted influential factors of rutting in a given year, one of which could be the weather.
-Rutting of at least $2$ millimetres at some locations (upper panel), or in a given year (lower panel), is concerning, given the expected lifetime of the pavement and the recommended levels for treatment.
-If maintenance is to be carried out at $20-25$ millimetres rut depth levels, depending on traffic, and pavement designs are expected to last between $20-30$ years, then some parts of the road would survive at most $12.5$ years ($10$ years in high-traffic areas).
-The minimum unexplained portion of rutting from $\hat{\omega}(s)$ is $-1.72$ mm/year.
-This suggests that those locations would have $1.72$ mm/year less rutting compared with mean ($0.01$ mm/year).
-That is, better real-life rutting performance than expected by the model.
-The rutting from the spatial component do appear to be relatively systematic, with a difference between the highest and lowest rutting values of $4.93$ millimetres.
-Including spatial components in the model allow us to explain more of the rutting variability and identify areas with high rutting rates not explained by known variables.
+4.1 Estimated spatial rutting component and yearly effect
+We analyze estimates from Model 1, identifying areas with high rutting rates not explained by AADT, pavement type, and lane width.
+The spatial rutting component ()$\hat{\omega}(s)$) plotted in Figure 3(a), varies significantly along the road.
+Some locations exhibit up to $3.15$ mm/year extra rutting beyond what is accounted for by known factors, while others show rutting exceeding $2$ mm/year.
+Conversely, the minimum unexplained rutting is $-2.45$ mm/year, indicating better performance than expected.
+Figure 3(b) presents estimates of the random yearly $\hat{\gamma}_{t}$.
+Rutting varies notably between years for spatial and non-spatial models, with minimum estimates in 2016 and maximum in 2018.
+A systematic shift between spatial and non-spatial models is evident, particularly in pavement type coefficients (see Table 3 in Section 4.2).
+Including spatial components in the model enhances rutting variability explanation and identifies areas with unexplained high rutting rates.
 This is confirmed in Figure 3 (b), where the rutting explained by the random yearly effect from the spatial model is at least $0.5$ mm/year less compared with the non-spatial model.
-We further investigate locations with unexpectedly high rutting rates in Section 4.2.
-Figure 3: Upper panel: Spatial rutting component (or unexplained portion of rutting $\{\hat{{\omega}}(s)\}$), with approximate 95% credible interval.
-Lower panel: Rutting from the random yearly effect ($\hat{{\gamma}_{t}}$) from Model 1 ( ) and Model 4 (       ), with approximate 95 % credible interval.
+Further discussion on estimated regression coefficients is in Section 4.2.
+Figure 3: Upper panel: Spatial rutting component (unexplained rutting, $\hat{{\omega}}(s)$), with 95% credible interval.
+Lower panel: Rutting from the random yearly effect ($\hat{{\gamma}_{t}}$) from Model 1 ( ) and Model 4 (       ), with 95 % credible interval.
+4.2 Estimated model parameters
+The estimated parameters and 95% credible intervals (CI) for spatial (Model 1) and non-spatial (Model 4) are shown in Table 3.
+Spatial models exhibit notably higher estimates for pavement type coefficients compared with non-spatial models (see also Table A.2 in Appendix A).
+Moreover, models lacking the spatial rutting component display larger variance in the ‘noise’ term ($\hat{\sigma}{\varepsilon}$), suggesting that the spatial rutting component captures unobservable spatially varying factors such as weather and traffic.
+Additionally, the spatial model better explains expected rutting, as indicated by smaller ‘noise’ terms ($\hat{\sigma}{\varepsilon}$) and uncertainty in random year effects ($\hat{\sigma}_{\gamma}$) compared with non-spatial models (refer to Table 3 and Table 4 in Appendix A).
+Further results are provided in Supplementary Materials LABEL:supplementary:addional_results.
+Table 3: Estimated model parameters with $95\%$ credible intervals
+Spatial Model Non-spatial Model
+Model 1 Model 4
+Mean/Median 95% CI Mean/Median 95% CI
+$\hat{\beta}_{\mathrm{Ac}}$ 7.09 [6.56, 7.61] 3.27 [2.84, 3.71]
+$\hat{\beta}_{\mathrm{Agc}}$ 7.31 [6.08, 7.73] 3.24 [2.78, 3.70]
+$\hat{\beta}_{\mathrm{Sma}}$ 3.53 [3.26, 3.80] 2.03 [1.84, 2.22]
+$\hat{\beta}_{\mathrm{d-1}}$ 0.76 [0.68, 0.84] 1.07 [0.98, 1.17]
+$\hat{\beta}_{\mathrm{w}}$ 0.07 [0.01, 0.13] 0.05 [-0.02, 0.11]
+$\hat{\sigma}_{\omega(s)}$ 0.48 [0.45, 0.52] - -
+$\hat{\sigma}_{\gamma}$ 0.92 [0.63, 1.42] 1.37 [0.94, 2.10]
+$\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.54 [1.53, 1.56]
+4.2.1 Effects of the explanatory variables
+Pavement types, particularly those with weaker binder concentrations like asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) and asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$), experience more rutting (Table 3).
+Qian et al.
+(2019) and Mascarenhas et al.
+(2020) shed light on how temperature and heavy traffic ($>$ 3.5 tonnes) impact rutting resistance on pavement types such as Agc and Ac.
+Notably, asphalt gravel concrete is predominantly situated approximately 30 kilometers from the Swedish border, where a sub-arctic climate, underdeveloped networks, and heavy traffic converge to adversely affect pavement performance Peel et al.
+(2007), Ketzler et al.
+(2021), Jenelius (2010), Hovi et al.
+(2018), Pinchasik et al.
+(2020).
+There is also greater uncertainty in asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), although considering spatial dependencies helps to explain more of the variability.
+Non-spatial models appear to explain variability between years rather than by pavement type, suggesting a nuanced relationship between these factors in understanding rutting susceptibility.
+Rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes more crucial without accounting for spatial variability (Model 4, Table 3).
+The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is minimal and positive, suggesting negligible impact on expected rutting.
+These findings indicate that changes in coefficients are significant when spatial dependencies of unmodeled factors (e.g., weather) are considered.
+Furthermore, the spatial model captures more uncertainty (smaller $\hat{\sigma}_{\varepsilon}$).
 4.3 Estimated rutting from the explanatory variables, random yearly effect and spatial rutting component
-We can also use the spatial model (Model 1) to identify locations, termed ‘hot spots’, that are in the greatest need of maintenance (Figure 4).
-For illustration, the model is used to infer rutting from the coefficients of the explanatory variables $\hat{\beta}$; the mean of the posterior means of the random yearly effects $\hat{\gamma}$; and the posterior mean of the spatial rutting component $\hat{\omega}(s)$.
-This is both the explained and unexplained rutting.
-The random yearly effect $\gamma_{t}$ varies from year to year, so we use an estimate of the mean, i.e., $\hat{\gamma}$.
-In addition, $\hat{\sigma}_{\xi_{t}}$ varies annually, but it is quite small (see Supplementary Materials LABEL:supplementary:addional_results) so it is not considered in our illustration.
-First, we estimate the rutting from the sum of the coefficients of the explanatory variables and the mean of the posterior means of the random yearly effects $\hat{\gamma}$ only.
-The results are from Model 1, so we denote this by: $\hat{\eta^{*}}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}$.
-Next, we estimate the rutting from the sum of the coefficients of the explanatory variables, the mean of the random yearly effect $\hat{\gamma}$, and posterior mean of the spatial rutting component $\hat{\omega}(s)$; we denote by $\hat{\eta}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}+\hat{\omega}(s)$.
-Since the different asphalt types; stone mastic asphalt (Sma), asphalt concrete (Ac) and asphalt gravel concrete (Agc), are laid at specific locations along E14, these are highlighted in various colours along the plot line $\hat{\eta^{*}}$.
-These results are shown in Figure 4.
-The effect of pavement type on rutting (i.e., the line showing the three asphalt types; Sma, Ac and Agc) is much less variable (ranging from $0.96$ to $3.89$ mm/year) compared with the sum of the explanatory variables, the mean of random yearly effect and the spatial rutting component (ranging from $0.03$ to $4.68$ mm/year); see graph (a).
-This suggests that the explanatory variables and the random yearly effect alone do not capture or realistically represent the variability in the expected rutting.
-Accounting for both spatial and random yearly effects allow more of the rutting variability to be explained.
-These can be seen in Table 3 and Table 4 in Appendix A, where $\hat{\sigma}_{\omega}(s)$ explains $0.47$ and $\hat{\gamma}$ explains $0.72$ compared with the remaining uncertainty ($\hat{\sigma}_{\varepsilon}=$ $1.38$) from the spatial model and the non-spatial version ($\hat{\sigma}_{\varepsilon}=$ 1.54).
-Additionally, the expected lifetime at some sections of the pavement is at most $6$ years considering only the explanatory variables and the random yearly effect or $5$ years considering the spatial rutting component $\hat{\omega}(s)$.
-This expected lifetime is depended upon traffic (i.e., $20$ or $25$ millimetres recommended maintenance for high- or low-traffic areas).
-We also see that the expected rutting $\hat{\eta}$ can be high, where the rutting is explained by the explanatory variables; or it could be moderate, but the spatial rutting component $\hat{\omega}(s)$ is large.
-Compared with the observed mean rutting across segments, which ranges from $-0.05$ to $8.33$ mm/year (see Figure 2, graph (a)), the maximum expected rutting estimated from the model (i.e., $4.68$ mm/year) is approximately halved the maximum of the observed mean rutting.
-In graph (b), the expected rutting from the sum of the explanatory variables, the random yearly effect and the spatial rutting component; lane width; and AADT (per ten thousand) are plotted for a short section of the road.
-This part of the road, from segments $1500-2700$ (or $30000-54000$ metres towards Sweden), is dominated by the weaker pavement type Agc and, as shown in graph (a), has spikes of accelerated rutting.
-When plotted against the AADT or lane width, we see that rutting is higher with increasing traffic and narrow road lanes.
-This large variability may, presumably, be a result of: (i) a change in asphalt type from asphalt concrete (Ac) to asphalt gravel concrete (Agc, see same road stretch in the graph (a)); (ii) the type of traffic in that area, e.g., heavy-traffic (vehicles $>3500$ kilograms); (iii) poor road design; or (iv) lack of timely and efficient pavement rehabilitation.
-The observed mean rutting (Figure 2 (a) in Section 2.4) is estimated to be $1.67$ mm/year.
-From graph (a), we found that the expected rutting, $\hat{\eta}$, for 351 road segments exceed this mean, and 22 of these road segments gave estimated rutting that exceeds $3$ mm/year (graphs (c) and (d), respectively).
-These 22 locations suggest potential hot spots for accelerated rutting and would require urgent maintenance as the expected lifetime would be at most 8 years, depending on traffic.
-(a) Expected rutting from Model 1. Dashed line is mean rutting $1.67$ mm/year.
-(b) Expected rutting, AADT and lane width
+We employ Model 1 to identify maintenance ‘hot spots’ where intervention is most critical.
+This model combines rutting inference from explanatory variable coefficients ${\hat{\beta}_{(\cdot)}}$, the mean of the yearly random effects ($\hat{\gamma}$), and the spatial rutting component ${\hat{\omega}(s)}$.
+Since the yearly random effects ($\gamma_{t}$) vary annually, we use an estimate of their mean, denoted as $\hat{\gamma}$.
+The yearly variation in ($\hat{\sigma}_{\xi_{t}}$) is small annually (Supplementary Materials LABEL:supplementary:addional_results), so it is not considered in our analysis.
+We account for both explained and unexplained rutting.
+Initially, rutting estimation relies on explanatory variables and the yearly random effect ($\hat{\eta}^{*}=\hat{\beta}_{{(1)}}z_{(1)(\mathrm{s,t})}+\hat{\gamma}$), with each asphalt type highlighted as Sma ( ), Ac ( ), and Agc ( ) (see Figure 4, graph (a)).
+Incorporating spatial rutting components yields a more comprehensive estimation ${\hat{\eta}=\hat{\beta}_{(1)}z_{(1)(\mathrm{s,t})}+\hat{\gamma}+\hat{\omega}(s)}$.
+A comparison of these estimates highlights the influence of spatial components.
+The results indicate less variability in rutting estimates without spatial components (ranging from $0.97$ to $3.64$ mm/year) compared to estimates including spatial factors (ranging from $-0.44$ to $4.67$ mm/year).
+This suggests that explanatory variables and the random yearly effect alone do not fully capture or represent the true rutting variability.
+Including spatial and yearly effects provides a more comprehensive understanding of rutting variability.
+Tables 3 and 4 in Appendix A demonstrate this, with the remaining uncertainty ($\hat{\sigma}_{\varepsilon}$) lower in the spatial model ($1.38$ mm/year) compared to the non-spatial model ($1.54$ mm/year).
+The findings indicate that high rutting estimates could arise from explanatory variables, while moderate estimates, coupled with large spatial rutting components, suggest the influence of other factors.
+(a) $\hat{\eta}^{*}$, $\hat{\eta}$ and mean rutting $1.62$ mm/year (       ).
+(b) $\hat{\eta}$, AADT and lane width
 Refer to caption (c) $\hat{\eta}>2$ mm Refer to caption (d) $\hat{\eta}>3$ mm.
-Figure 4: Graph (a): Expected rutting from: i) the spatial rutting component $\hat{\omega}(s)$, the random yearly effects plus coefficients of all explanatory variables, interacting with AADT for Model 1 $\left\{\hat{\eta}\right\}$ and ii) from the sum of the explanatory variables and the of random yearly effect only $\left\{\hat{\eta^{*}}\right\}$, with each asphalt type highlighted.
+Figure 4: Graph (a): Expected rutting from: i) $\hat{\eta}$ and ii) from $*\hat{\eta}{{}^{*}}$, with each asphalt type highlighted.
 The effects are inferred for the year 2020.
-Graph (b): Expected rutting plotted against ADDT and lane width for a small stretch of E14 (closer to Sweden).
-Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year or $3$ mm/year, respectively.
-4.2.1 Investigating areas with high rutting rates
-The large spike in Figure (4) (a), showing high rutting rates, is further investigated.
-We look at the road segments 2100 - 2500 metres surrounding this spike.
-This area is the municipality of Meråker, just 20 kilometres west of Storlien.
-This stretch, as shown in Figure 4 (a), gave high expected rutting, reaching $4.68$ millimetres, with an expected lifetime of just $5$ years (low-traffic area with required maintenance at 25 millimetres rut depth).
-This accelerated rutting could be because of clogged drainage, flooding, missing side ditches (Figure 5), heavy-traffic or road designs with railguards installed close to the lane edge, allowing that traffic drives more within the same wheel tracks.
-Meråker has industry and agriculture, therefore, roadways are more subjected to heavy vehicles, including farm vehicles such as tractors (industry, wheel drive, row crop).
-(a) Short section between EV14 S6D1 m6148-m8212 F1 with high rutting.
-(b) Location EV14 S6D1 m8017 F1 in year 2020
-(c) Location EV14 S6D1 m7971 F1 in year 2021
-(d) Location EV14 S6D1 m7961 F1 in year 2022
-Figure 5: Graph (a): Location with high rutting rates (road segments $2100-2500$ towards Sweden).
-Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or go to the next url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@337286,7036109,14/vegsystemreferanse:337312.447:7036064.784
-To further investigate this road section with high rutting values, we consider the effect of the explanatory variables, random yearly effect, and the spatial rutting component.
-We explore the spatial Model 1 using various values of the explanatory variables to approximate the rutting.
-First, we use the true values in the data, i.e., the asphalt type, the lane width and rut depth measured at that location, denoted by $\hat{\eta}_{\mathrm{True}}$.
-Second, the asphalt type is changed from its true state of asphalt gravel concrete (Agc) to asphalt concrete (Ac) or stone mastic asphalt (Sma); represented by $\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$, respectively.
-Third, the ‘best’ (or ‘worse’) values are used, i.e., if pavement type with the strongest binder, Sma, or a rut depth value from the previous year of $0.01$ millimetres ($\sim 0$), or lane width of $6$ metres were used.
-This is denoted by $\hat{\eta}_{\mathrm{Best}}$.
-These values are chosen following the reasoning: ($\sim$) zero rut depth means no rutting; wider roads mean a slower rate of rutting.
-The values considered to be ‘worse’, resulting in accelerated rutting, include a rut depth from the previous year of 25 millimetres, lane width of 2.75 metres and asphalt gravel concrete (Agc) pavement type.
-This is denoted by $\hat{\eta}_{\mathrm{Worse}}$.
-These values are found in the E14 data set.
-The cases are plotted in Figure 6 (a), which gives the expected rutting from the sum of the explanatory variables, th yearly random effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$.
-Using the worse values in the analysis, the expected rutting is at least 2.3 ($2.25-7.94$) mm/year; Figure 6, graph (a) ( ).
-This means that with low AADT and the recommended maintenance rut depth level of 25 millimetres, pavement deterioration would accelerate and treatment or new pavement would be required within $3$ to $11$ years after initial construction.
-The best covariate values ( ) gave at most $4.26$ mm/year of rutting, with the least rutting value ($0.21$ mm/year).
-And, the earliest maintenance activity would be required from around 6 years after construction or maintenance.
-Further, the best values ( ) gave similar estimates of annual rutting to that of using the stronger asphalt binder type Sma (stone mastic asphalt,  ), except at road segments $2279-2388$, where the estimated rutting from only changing the asphalt type from asphalt gravel concrete to stone mastic is lower than that of using the best values.
-This suggests that asphalt type is the driving factor of expected rutting.
-Pavement types asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) gave similar estimates of expected rutting.
-This result is expected because they are made with similar binder concentration strengths.
+Graph (b): Expected rutting plotted against AADT and lane width for a short stretch.
+Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year and $3$ mm/year, respectively.
+4.3.1 Comparing expected rutting with observed data
+The observed mean rutting (Figure 2 (a) in Section 2.4) stands at $1.62$ mm/year.
+In Figure 4, graph (a), 361 road segments have an expected rutting exceeding this mean, with 49 segments showing estimated rutting exceeding $3$ mm/year (graphs (c) and (d), respectively).
+These 49 locations signify potential hot spots for accelerated rutting, necessitating urgent maintenance, as their expected lifetime could be as short as 8 years, depending on traffic.
+Additionally, the observed mean rutting across segments ranges from -0.14 to 8.33 mm/year (Figure 2, graph (a)), while the maximum expected rutting from the model is 4.67 mm/year (Figure 4, graph (a)), approximately half the maximum observed mean rutting.
+In Figure 4, graph (b), we plot the expected rutting $\hat{\eta}$, lane width, and AADT (per ten thousand) for a short road section spanning segments $1500-2700$ (or $30-54$ km from Stjørdal), characterized by weaker pavement type Agc.
+This section exhibits spikes of accelerated rutting, particularly evident with increasing traffic.
+Possible factors contributing to this variability include changes in asphalt type, traffic composition, suboptimal road design, or inadequate pavement maintenance.
+4.4 Investigating an area with high rutting rates
+The significant spike in Figure (4) (a), indicating high rutting rates, prompts further investigation.
+We examine road segments spanning 2100 to 2700 meters surrounding this spike, located in the municipality of Meråker, just 20 kilometers west of the Swedish border at Storlien.
+This stretch, depicted in Figure 4 (a), exhibits notably high rutting, peaking at $4.67$ mm, with an estimated lifespan of only $5$ years (in a low-traffic area requiring maintenance at 25 mm rut depth).
+Possible contributors to this accelerated rutting include clogged drainage, flooding, absence of side ditches (see Figure 5), heavy traffic, or road designs featuring railguards installed close to the lane edge, which encourages traffic to remain within the same wheel tracks (refer to Figure A.4 in Appendix A.4).
+Meråker’s industrial and agricultural activities result in increased traffic from heavy vehicles, including farm equipment such as tractors (industrial, wheel drive, row crop).
+(a) Short section between EV14 S6D1 m3841-m4285 F1 with high rutting.
+(b) Location EV14 S6D1 m4068 F1 in year 2020
+(c) Location EV14 S6D1 m4059 F1 in year 2021
+(d) Location EV14 S6D1 m4065 F1 in year 2022
+Figure 5: Graph (a): Location with high rutting rates (road segments $2100-2700$ towards Sweden).
+Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@335074,7038866,15/vegsystemreferanse:335035.067:7038844.352
+4.4.1 Exploring the spatial model with various values of the explanatory variables
+To further investigate high rutting values in a road section, we analyze the impact of explanatory variables, yearly variation, and spatial rutting component ($\hat{\eta}$).
+We explore spatial Model 1 using different values for variables to estimate rutting.
+First, we use actual measurements like asphalt type, lane width, and rut depth ($\hat{\eta}_{\mathrm{True}}$).
+Second, we consider alternative asphalt types - asphalt concrete (Ac) or stone mastic asphalt (Sma) ($\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$).
+Third, we examine extreme values such as the strongest pavement (Sma), minimal rut depth (0.01 mm), or wider lanes (6 meters) ($\hat{\eta}_{\mathrm{Best}}$).
+Conversely, accelerated rutting scenarios involve deeper rut depth (25 mm), narrower lanes (2.75 meters), and asphalt gravel concrete (Agc) ($\hat{\eta}_{\mathrm{Worse}}$) from the E14 dataset.
+These scenarios are depicted in Figure 6 (a).
+With worse values ( ), rutting is projected to be at least 1.5 mm/year ($1.76-8.20$), indicating accelerated deterioration and the need for maintenance within 3 to 14 years.
+Locations with unexpectedly high rutting, exceeding 4 mm/year, are observed, possibly due to increased AADT compared to other areas (Figure 4, graph b).
+Using the best covariate values ( ), rutting is projected to be at most 4.30 mm/year, with maintenance required approximately 6 years after construction or maintenance.
+Additionally, using the stronger asphalt binder Sma yields similar rutting estimates but generally lower than those using the best values, indicating asphalt type as a primary factor.
+Asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) show similar rutting estimates due to similar binder concentrations.
 (a) Expected rutting at road segments $2100-2500$ (metres) towards Sweden.
@@
 (e) SCI/BCI for the whole E14.
-Figure 6: Graph (a): Expected rutting inferred from the spatial rutting component $\hat{\omega}(s)$ plus the ‘best’ (or ‘worse’) explanatory values at road segments $2100-2500$ towards Sweden.
-The estimate $\hat{\eta}_{\mathrm{Agc}}$ is for pavement type Agc plus the random yearly effect and the spatial rutting component.
-Graph (b): expected rutting inferred from $\hat{\omega}(s)$ plus the random yearly effect $\hat{\gamma}$ and each explanatory variable added sequentially in the model.
-Graph (c): bearing capacity for year 2021 (tonn), with 16 tonn or greater indicating a good parameter.
-Graphs (d) and (e) give the ratios of surface curvature index (SCI) to base curvature index (BCI) for road segments 2100-2500 (metres) or the whole E14, respectively, with values below $12$ indicating a good curvature index.
-The above results suggest pavement type as the driving explanatory variable for expected rutting (Figure 6), so we determine its importance on rutting by changing the asphalt type from the true one (i.e., asphalt gravel concrete, Agc) to the stronger asphalt type Sma (stone mastic asphalt) and keeping all other variables (e.g., lane width) as given in the data set.
-The remaining explanatory variables are then added sequentially in the model with pavement type, i.e., stacking.
-Pavement type (Agc or Sma) is first added to the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$.
-We denote this by: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$, respectively.
-Next, rut depth from the previous year is added, and it is represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$.
-Similarly, the model with lane width $\{\hat{\beta}_{w}z_{\mathrm{w(s,t)}}\}$ added is represented by $\hat{\tilde{\eta}}_{\mathrm{w}}$.
-The results are shown in Figure 6, graph (b).
-Results from using stone mastic asphalt (Sma) are represented by the solid lines, and from using asphalt gravel concrete (Agc), the dotted lines.
-The results confirm that pavement type is the driving factor of rutting.
-If the stronger binder type Sma were laid at this stretch instead of the true pavement type Agc, expected rutting would decrease, reaching at most $4.35$ mm/year (compared with 4.68 mm/year).
-This is approximately 0.33 mm/year less rutting annually, with an expected lifetime of 6 years, i.e., 1 extra year.
-We also look at the available data for bearing capacity (Figure 6, graph (c)), surface curvature index (SCI) and base curvature index (BCI) for road segments 2100-2500.
-The data is available for year 2021, but it is a good indicator of the construction quality and the soil to sustain pressure over the past decade; Figure 6, graph (c).
-For bearing capacity, values greater than $16$ tonns (       ) suggest a good foundation design.
-The ratio of SCI to BCI ((Figure 6, graphs (d) and (e)) gives an indication of the location of the deflection (deformation) in the pavement layer (surface or base).
-The data suggests good bearing capacity, and that high rutting may be a result of other influential factors, e.g., side slope height (data not available), lack of side ditches to effect water drainage or deformation in the sub-base layer developed in thawing periods.
-We note that the bearing capacity data was sampled during the summer and corrected for temperature.
-During the summer, the measurements can be higher due to dry conditions, compared to those in thawing periods with high moisture content.
-Permanent deformation could have occurred during thawing periods, resulting in high rutting.
-There is also the possibility of varying construction material quality at various parts of the road, even though the pavement type (asphalt gravel concrete) is the same for the entire stretch (road segments $1500-3355$).
-Therefore, we investigate the consistency in the pavement construction by looking at deflections in the surface layer and the subsoil or reinforcement layer.
-The ratio of SCI to BCI is a good indication of the strengths or weaknesses in the construction, with $\mathrm{SCI/BCI}<12$ (       ) indicating a good parameter value.
-If this ratio is constant, then the construction is considered even, otherwise, there are variations indicating deformation in the various layers.
-The data, suggests an uneven construction of E14, with deeper deflections in the base layer.
-We conclude that pavement type is the most important contributing factor to rutting. So, stronger binder types are recommended to ensure longer expected lifetimes.
-4.3 Estimated model parameters
-The estimated parameters and 95% credible intervals (CI) for the spatial Model 1 and non-spatial Model 4 are reported in Table 3.
-Comparing the coefficients for the explanatory variables between spatial models and their non-spatial version, we observe that the spatial models have considerably larger estimates for the regression coefficients of pavement type (see also Figure A.2 in Appendix A).
-Further, we see that the variance of the ‘noise’ term ($\hat{\sigma}_{\varepsilon}$) is larger for the models without the spatial rutting component.
-One explanation for this result is that the spatial rutting component account for unobservable factors, e.g., weather that vary in space, the effect of studded tires and other influencing factors such as traffic type that are not included in the model.
-Varying weather patterns across Norway and towards the Swedish border will affect the function of surface layers of pavement construction.
-In the winter time, increased temperatures result in a shorter time of frozen ground and snow on the surface and changes in the number of freeze-thaw cycles, thus, negatively affecting lower-grade binder types such as asphalt gravel concrete (Agc) and leading to increasing rutting.
-Asphalt gravel concrete is laid from () $(\sim)\ 30$ kilometres towards the Swedish border (Figure 1 (a)); an area that lie in the sub-artic climate zone with long severe winters and short warm to cool summers Peel et al.
-(2007), Ketzler et al.
-(2021), and having poorly developed networks (rural E14 highway west of Östersund) Jenelius (2010).
-This area is also more likely to be traversed by heavy-traffic (exceeding $3500$ kilograms or $3.5$ tonnes) Hovi et al.
-(2018), Pinchasik et al.
-(2020), negatively affecting pavement performance.
-More of the expected rutting is also explained by rut depth and lane width from the spatial model compared with the non-spatial model.
-This can be seen by the smaller values of the ‘noise‘ terms ($\hat{\sigma}_{\varepsilon}$) and uncertainty in the random year effect ($\hat{\sigma}_{\gamma}$) from the spatial model (see Table 3 and Table 4 in Appendix A for the simpler models).
-Table 3: Estimated model parameters with $95\%$ credible intervals
-Spatial Model Non-spatial Model
-Model 1 Model 4
-Mean/Median 95% CI Mean/Median 95% CI
-$\hat{\beta}_{\mathrm{Ac}}$ 7.13 [6.62, 7.65] 3.25 [2.81, 3.68]
-$\hat{\beta}_{\mathrm{Agc}}$ 7.34 [6.92, 7.76] 3.21 [2.75, 3.67]
-$\hat{\beta}_{\mathrm{Sma}}$ 3.63 [3.37, 3.90] 2.00 [1.81, 2.19]
-$\hat{\beta}_{\mathrm{d-1}}$ 0.74 [0.66, 0.82] 1.08 [0.98, 1.18]
-$\hat{\beta}_{\mathrm{w}}$0.10 [-0.14, -0.07] 0.05 [-0.02, 0.11]
-$\hat{\sigma}_{\omega(s)}$ 0.47 [0.44, 0.51] - -
-$\hat{\sigma}_{\gamma}$ 0.72 [0.62, 0.88] 1.38 [0.96, 2.16]
-$\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.54 [1.53, 1.56]
-4.3.1 Effects of the explanatory variables
-Interacting with the AADT, which is spatially varying, the effect of asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) on rutting may be explained by: heavy-traffic ($>3.5$ tonnes), whose proportion changes along the road and poor network design as these are constant in time.
-Also, various parts of the pavement have several layers due to non-removal during maintenance, leading to variability in pavement performance.
-The uncertainty from asphalt concrete and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) are also larger (wider 95% credible interval) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), but accounting for spatial dependencies allows more of the variance to be explained.
-Rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes more important when the spatial variability in pavement rutting is not considered (Model 4, Table 3).
-Further, instead of explaining the variability in expected rutting by pavement type, the non-spatial models explain this variability between years (Table 3 and Figure 3 (b)).
-The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is small and negative.
-This suggests that narrower lane widths would give higher expected rutting.
-Based on these results, we conclude that changes in the coefficients are large when spatial dependencies of factors that are not modelled because of non-observable data (e.g., weather) are considered.
-And, the uncertainty explained is greater (smaller $\hat{\sigma}_{\varepsilon}$) for the spatial model.
+Figure 6: Graph (a) displays expected rutting using spatial rutting component $\hat{\omega}(s)$ for road segments $2100-2700$, along with the ’best’ or ’worst’ explanatory values.
+$\hat{\eta}_{\mathrm{Agc}}$ represents pavement type Agc, random yearly effect $\hat{\gamma}$, and $\hat{\omega}(s)$.
+Graph (b) illustrates expected rutting inferred from $\hat{\omega}(s)$, $\hat{\gamma}$, and sequentially added explanatory variables.
+Graph (c) shows 2021 bearing capacity in tons; values $\geq$ 16 indicate good parameters.
+Graphs (d) and (e) depict SCI to BCI ratios for a short stretch or the entire E14; values $<$ 12 indicate good curvature index.
+We examine the impact of pavement type on rutting (Figure 6, graph (b)).
+Changing from asphalt type Agc (asphalt gravel concrete) to the stronger Sma (stone mastic asphalt) while keeping all other variables constant, we sequentially add additional explanatory variables to the model, starting with pavement type.
+Initially, pavement type (Agc or Sma) is included with the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$ as follows: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$.
+Subsequently, the previous year’s rut depth is incorporated, represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$.
+Similarly, the model with added lane width ${\hat{\beta}_{w}z_{\mathrm{w(s,t)}}}$ is denoted as $\hat{\tilde{\eta}}_{\mathrm{w}}$.
+Results from using stone mastic asphalt (Sma) are depicted by solid lines, and asphalt gravel concrete (Agc) results are shown with dotted lines.
+The findings affirm that pavement type significantly influences rutting.
+Implementing the stronger binder type Sma instead of the actual pavement type Agc would decrease expected rutting, reaching a maximum of $4.35$ mm/year (compared to 4.68 mm/year), approximately reducing rutting by 0.33 mm/year annually.
+Over a projected lifetime of 6 years, this equates to an additional year of serviceability.
+4.4.2 Evaluating pavement subgrade support on expected rutting
+We examine bearing capacity, surface curvature index (SCI), and base curvature index (BCI) for road segments 2100-2700 using 2021 data, indicative of construction quality and soil pressure sustainability over the past decade (Figure 6, graph (c)).
+Bearing capacity values exceeding 16 tons (       ) suggest a robust foundation design.
+The SCI to BCI ratio (Figure 6, graphs (d) and (e)) reveals deflection location within the pavement layers.
+High rutting may result from factors like side slope height (data unavailable), inadequate drainage, or sub-base layer deformation during thawing periods.
+Bearing capacity data, sampled and temperature-corrected during summer, may inflate measurements compared to thawing periods.
+Varying construction material quality along the road stretch (segments 1500-3355) warrants investigation into surface and subsoil layer deflections.
+An SCI/BCI ratio below 12 (       ) indicates robust construction.
+Inconsistent ratios suggest layer deformation, with data indicating uneven construction on E14, featuring deeper base layer deflections.
+In conclusion, pavement type significantly influences rutting, recommending stronger binder types for extended lifespan.
 5 How road authorities can make use of this model
-The spatial-statistical model serves as a potent tool for road authorities, offering a data-driven approach to enhance maintenance strategies, identify critical areas for intervention and optimize road infrastructure durability.
-This approach facilitates a more ultimate diagnosis of rutting, empowering authorities to make informed decisions for sustainable road infrastructure.
-Road authorities can use the spatial-statistical model for historical rutting data in several ways to improve maintenance strategies and prioritize interventions.
-These include; i) the identification of ‘hot spots’ — the spatial model helps pinpoint hot spots where rutting is accelerating, indicating areas in urgent need of maintenance.
-By analyzing coefficients of factors known to influence rutting (e.g., asphalt type), random effects, and spatial components, authorities can prioritize resources for targeted interventions in these critical locations; ii) comprehensive rutting diagnostic and understanding — the model provides a more comprehensive diagnostic and understanding of rutting.
-Explanatory variables and random yearly effects alone may not capture the full variability, but the inclusion of spatial components enhances the model’s explanatory power.
-This comprehensive understanding allows road authorities to make informed decisions about maintenance priorities; iii) extended expected lifetime — considering both spatial and random yearly effects extends the expected lifetime estimation of road sections.
-This information is valuable for long-term planning, enabling authorities to proactively address maintenance needs and allocate resources efficiently; iv) pavement type considerations — the model highlights the significant impact of pavement type on rutting rates.
-Road authorities can use this information to make informed decisions about pavement materials during construction or rehabilitation, optimizing durability and extending the overall lifetime of the road infrastructure.
-The recommendation of stronger asphalt binders, such as stone mastic asphalt (Sma), to enhance pavement durability, provides practical guidance for selecting materials that contribute to longer-lasting road surfaces; v) investigation of contributing factors — in-depth investigations into high rutting areas reveal potential contributing factors beyond pavement type, such as drainage issues, heavy traffic, and road design flaws.
-Road authorities can address these factors to prevent or mitigate rutting, ensuring a more holistic approach to road maintenance; vi) spatial model advantages — comparisons with non-spatial models highlight the advantages of including spatial components.
-The larger estimates for pavement type coefficients in spatial models emphasize the importance of spatial factors in explaining rutting variations.
-This information reinforces the importance of using spatial models for accurate and reliable inference and predictions; and finally vii) data-driven decision-making — the estimated model parameters and credible intervals provide a robust foundation for data-driven decision-making.
-Road authorities can rely on these parameters to make informed choices regarding maintenance priorities, resource allocation, and infrastructure improvements.
+The spatial-statistical model serves as a potent tool for road authorities, offering a data-driven approach to enhance maintenance strategies, identify critical areas for intervention, and optimize road infrastructure durability.
+This approach facilitates a more precise diagnosis of rutting, enabling authorities to make informed decisions for sustainable road infrastructure.
+Our analyses reveal that rutting exceeding 2 mm/year is expected at several locations along E14.
+Given the expected lifetime of the pavement and recommended treatment levels, this level of rutting is concerning.
+Maintenance carried out at rut depths of 20-25 mm, depending on traffic, and pavement designs expected to last between 20-30 years indicate that some sections of the road may only survive for at most 12.5 years (10 years in high-traffic areas).
+It’s imperative to consider spatial rutting components; models fitted with variables known to influence rutting and random yearly effects alone underestimate rutting variability, thus inflating pavement lifetime estimates.
+Including the spatial rutting component reduces estimated pavement lifetimes to at most 6 years, compared to 5 years without it.
+This information is crucial for road authorities’ long-term planning and resource allocation.
+We’ve demonstrated the spatial-statistical model’s ability to identify areas with unexpectedly high rutting, pinpointing hot spots where rutting is accelerating and indicating urgent maintenance needs.
+By analyzing factors influencing rutting, random effects, and spatial components, authorities can prioritize resources for targeted interventions.
+Additionally, the model provides a comprehensive understanding of rutting, enhancing authorities’ ability to make informed decisions about maintenance priorities and pavement materials during construction or rehabilitation.
+The model can investigate contributing factors to rutting beyond pavement type, such as drainage issues, heavy traffic, and road design flaws.
+Comparisons with non-spatial models emphasize the importance of spatial factors in explaining rutting variations, reinforcing the need for spatial models for accurate inference and predictions.
+The estimated model parameters and credible intervals provide a robust foundation for data-driven decision-making, enabling road authorities to make informed choices regarding maintenance priorities and infrastructure improvements.
 6 Conclusion
-This research provides valuable insights into the factors influencing pavement rutting and the measurement of rut depth.
-The study confirms that the selection of asphalt type significantly affects rutting performance, with rut-resistant mixtures proving to be more effective in mitigating rutting.
-Lane width also plays a role in pavement rutting.
-Narrower roads tend to experience more concentrated traffic loads, which can lead to increased rutting.
-Wider roads, on the other hand, distribute the traffic load over a larger area, reducing the likelihood of rut formation.
-The data analysis indicated that narrower road stretches experienced higher rut depths.
-This outcome can be attributed to increased traffic volumes and higher wheel loads on roads, leading to more significant rutting.
-The findings, therefore, stress the importance of considering lane width when designing and maintaining pavements to minimize rutting distress.
-The recommendation is to incorporate a life-cycle cost-benefit perspective in the road design and planning process to minimize the potential for rutting issues over time.
-Further research is encouraged to explore additional factors influencing rutting; some of which include bearing capacity (Hopkins et al.
-1998) as inadequate bearing strength of soil subgrades will not withstand construction traffic loading; ditch depth information that allows outflow of water from pavement structure; and lane width.
-This would likely increase the credibility of research results and the applicability of models to other road networks.
-In addition, it would be worthwhile to evaluate the long-term performance of rut-resistant asphalt mixtures in various traffic and environmental conditions.
+This research offers valuable insights into the rutting phenomenon by elucidating both well-established factors and those that remain unexplained.
+It shows the importance of incorporating spatial information to accurately estimate rutting variability, showing that observations in close proximity exhibit higher similarity than those at a distance.
+To achieve a comprehensive understanding, it is imperative to consider both observable elements like traffic and temperature, as well as unobservable factors like weather, and integrate spatial rutting components.
+Moreover, the study confirms the significant impact of asphalt type selection on rutting performance, with rut-resistant mixtures proving more effective in mitigating rutting.
+Lane width also influences pavement rutting, with narrower roads experiencing more concentrated traffic loads and, consequently, increased rutting.
+Conversely, wider roads distribute traffic loads over a larger area, reducing rut formation.
+The analysis reveals that wider road sections tend to experience higher rut depths compared to narrower road stretches.
+This result can be attributed to the increased traffic volumes and higher wheel loads present on wider roads, which contribute to more pronounced rutting.
+To address this issue and minimize rutting problems over time, it is advisable to integrate a life-cycle cost-benefit perspective into road design and planning processes.
+Further research is encouraged to explore additional factors influencing rutting, such as bearing capacity, ditch depth for water outflow, and lane width.
+This would enhance the credibility of research results and the applicability of models to other road networks.
+Additionally, evaluating the long-term performance of rut-resistant asphalt mixtures under various traffic and environmental conditions would be worthwhile.
 Acknowledgement
@@
 Funding
-This work was supported by the Norwegian University of Science and Technology in collaboration with Statens Vegvesen Project [grant number 9056608 – SMARTere Vedlikehold (2019-2023)].
+This work was supported by the Norwegian University of Science and Technology in collaboration with Statens Vegvesen Project [grant number 9056608 – Smarter Maintenance (2019-2023)].
 Appendix A Inference from model parameters
@@
 Table 4: Estimated model parameters with $95\%$ credible intervals
-SPATIAL MODELS Model 1 Model 2 Model 3 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 7.13 [6.62, 7.65] 7.19 [6.66, 7.72] 6.98 [6.44, 7.53] $\hat{\beta}_{\mathrm{Agc}}$ 7.34 [6.92, 7.76] 7.38 [6.95, 7.81] 7.15 [6.72, 7.59] $\hat{\beta}_{\mathrm{Sma}}$ 3.63 [3.37, 3.90] 3.57 [3.30 3.85] 3.36 [3.08, 3.64] $\hat{\beta}_{\mathrm{d-1}}$ 0.74 [0.66, 0.82] 0.77 [0.69, 0.85] - - $\hat{\beta}_{\mathrm{w}}$0.10 [-0.14, -0.07] - - - $\hat{\sigma}_{\omega(s)}$ 0.47 [0.44, 0.51] 0.48 [0.45, 0.53] 0.52 [0.48, 0.57] $\hat{\sigma}_{\gamma}$ 0.72 [0.62, 0.88] 0.81 [0.64, 1.09] 0.72 [0.58, 0.94] $\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.38 [1.37, 1.39] 1.39 [1.38, 1.40] NON-SPATIAL MODELS Model 4 Model 5 Model 6 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 3.25 [2.81, 3.68] 3.25 [2.81, 3.69] 2.72 [2.28, 3.16] $\hat{\beta}_{\mathrm{Agc}}$ 3.21 [2.75, 3.67] 3.20 [2.74, 3.67] 2.71 [2.24, 3.18] $\hat{\beta}_{\mathrm{Sma}}$ 2.00 [1.81, 2.19] 2.02 [1.83, 2.21] 1.55 [1.36, 1.73] $\hat{\beta}_{\mathrm{d-1}}$ 1.08 [0.98, 1.18] 1.08 [0.98, 1.18] - - $\hat{\beta}_{\mathrm{w}}$ 0.05 [-0.02, 0.11] - - - - $\hat{\sigma}_{\gamma}$ 1.38 [0.96, 2.16] 1.45 [0.98, 2.21] 1.43 [0.99, 2.25] $\hat{\sigma}_{\varepsilon}$ 1.54 [1.53, 1.56] 1.54 [1.53, 1.56] 1.56 [1.54, 1.57]
+SPATIAL MODELS Model 1 Model 2 Model 3 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 7.09 [6.56, 7.61] 7.08 [6.55, 7.61] 6.89 [6.34, 7.44] $\hat{\beta}_{\mathrm{Agc}}$ 7.31 [6.88, 7.73] 7.29 [6.86, 7.71] 7.07 [6.64, 7.50] $\hat{\beta}_{\mathrm{Sma}}$ 3.53 [3.26, 3.80] 3.55 [3.28 3.82] 3.31 [3.03, 3.59] $\hat{\beta}_{\mathrm{d-1}}$ 0.76 [0.68, 0.84] 0.76 [0.68, 0.84] - - $\hat{\beta}_{\mathrm{w}}$ 0.07 [0.01, 0.13] - - - $\hat{\sigma}_{\omega(s)}$ 0.48 [0.45, 0.52] 0.48 [0.45, 0.52] 0.52 [0.48, 0.57] $\hat{\sigma}_{\gamma}$ 0.92 [0.63, 1.42] 0.93 [0.63, 1.40] 0.93 [0.63, 1.42] $\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.38 [1.38, 1.39] 1.39 [1.38, 1.40] NON-SPATIAL MODELS Model 4 Model 5 Model 6 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 3.27 [2.84, 3.71] 3.28 [2.85, 3.72] 2.72 [2.29, 3.16] $\hat{\beta}_{\mathrm{Agc}}$ 3.24 [2.78, 3.70] 3.23 [2.77, 3.70] 2.71 [2.25, 3.17] $\hat{\beta}_{\mathrm{Sma}}$ 2.03 [1.84, 2.22] 2.05 [1.86, 2.24] 1.55 [1.36, 1.73] $\hat{\beta}_{\mathrm{d-1}}$ 1.07 [0.98, 1.17] 1.08 [0.98, 1.18] - - $\hat{\beta}_{\mathrm{w}}$ 0.05 [ -0.02, 0.11] - - - - $\hat{\sigma}_{\gamma}$ 1.37 [0.94, 2.10] 1.37 [0.94, 2.10] 1.41 [0.96, 2.17] $\hat{\sigma}_{\varepsilon}$ 1.54 [1.53, 1.56] 1.56 [1.54, 1.57] 1.56 [1.54, 1.57]
 A.1 Model Comparison
@@
 DIC WAIC (-) loglikelihood DIC WAIC (-) loglikelihood
-1 193678.05 287136.55 -101455.25 4 104290.62 100995.92 -52229.45
-2 194145.55 301922.59 -101459.74 5 104292.58 101057.74 -52223.40
-3 193938.88 285335.82 -101625.45 6 104763.18 101615.56 -52453.58
+1 195067.76 195660.91 -101464.72 4 104290.61 100996.40 -52229.53
+2 195154.12 195669.17 -101460.32 5 104290.62 100996.45 -52223.62
+3 195242.58 195959.89 -101624.69 6 104763.10 101556.00 -52453.66
 A.2 Marginal Posterior Distributions of Explanatory Variables
@@
 Narrower densities suggest smaller estimated uncertainty.
-We find that some of the uncertainty of the Gaussian observations $(\hat{\sigma}_{\varepsilon})$ and random yearly effect $(\hat{\sigma}_{\gamma})$ go into the spatial effect $\{\hat{\sigma}_{\omega(s)}\}$, and this can be interpreted as some of the unobserved variability (e.g., weather, temperature) being explained by the spatial component; see Table 3 and Figure A.3 (a) and (b).
+We find that some of the uncertainty of the Gaussian observations $(\hat{\sigma}_{\varepsilon})$ and random yearly effect $(\hat{\sigma}_{\gamma})$ go into the spatial effect $\{\hat{\sigma}_{\omega(s)}\}$, and this can be interpreted as some of the unobserved variability (e.g., weather, temperature) being explained by the spatial rutting component; see Table 3 and Figure A.3 (a) and (b).
 This is confirmed by the smaller estimates of the ‘noise’ term $(\hat{\sigma}_{\varepsilon})$ compared with the non-spatial model (Table 4).
 We examine the effects of the range of the common spatial field ($\hat{\rho}$), and the standard deviation of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$.
-As more explanatory variables are added to model (e.g., rut depth), the variability of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$ decreases; see Figure A.3, graph (d).
+As more explanatory variables are added to model (e.g., rut depth), the variability of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$ decreases; see Table 4 and Figure A.3, graph (d).
 In Model 2, rut depth from the previous year is included, and this effect is large with the credible interval not covering zero.
@@
 Figure A.3: Estimates of marginal posterior distributions of the spatial models (solid lines) and non-spatial models (dotted lines).
+A.4 Area with high rutting rates
+Figure A.4 shows the location where rutting exceeds $7$ mm when the worse values of the explanatory variables are used to explore Model 1 (see Section 4.4)
+(a) Short section between EV14 S6D1 m6802-m6882 F1 with high rutting.
+(b) Location EV14 S6D1 m6881 F1 in year 2020
+(c) Location EV14 S6D1 m6808 F1 in year 2021
+(d) Location EV14 S6D1 m6807 F1 in year 2022
+Figure A.4: Graph (a): Location with high rutting rates (road segments $2100-2700$ towards Sweden).
+Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or go to the next url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@337039,7037105,18/vegsystemreferanse:337039.29:7037149.25

### v3

@@
-A Spatial-statistical model to analyse historical rutting data
+interacttfvsample_Supplementary
 Abstract
@@
 Our model provides a thorough diagnostic analysis, pinpointing areas exhibiting unexpectedly high rutting rates.
-Incorporating spatial and random components, and important explanatory variables like annual average daily traffic (traffic intensity), pavement type, rut depth and lane width, our proposed models account for and estimate the influence of these variables on rutting.
+Incorporating spatial and random components, and important explanatory variables like annual average daily traffic (traffic intensity), asphalt type, rut depth and lane width, our proposed models account for and estimate the influence of these variables on rutting.
 This approach not only quantifies uncertainties and discerns locations at the highest risk of requiring maintenance, but also uncover spatial dependencies in rutting (millimetre/year).
 We apply our models to a data set spanning eleven years (2010-2020).
-Our findings emphasize the systematic unexplained spatial rutting effect, where some of the rutting variability is accounted for by spatial components.
-Pavement type, in conjunction with traffic intensity, is also found to be the primary driver of rutting.
+Our findings emphasise the systematic unexplained spatial rutting effect, where some of the rutting variability is accounted for by spatial components, asphalt type, in conjunction with traffic intensity, is also found to be the primary driver of rutting.
 Furthermore, the spatial dependencies uncovered reveal road sections experiencing more than 1 millimeter of rutting beyond annual expectations.
@@
 Flexible pavement; predictive maintenance; rutting; spatial-statistical model; uncertainty
-articletype: ARTICLE TEMPLATE
 1 Introduction
@@
 2023) and these models typically include rutting.
-Empirical models usually relate the accumulation of distresses (e.g., rut, roughness, cracks) in the pavement layers to the load repetitions (annual average daily traffic) and other affecting factors (e.g., pavement or asphalt type, weather, width of the road, layer thickness).
+Empirical models usually relate the accumulation of distresses (e.g., rut, roughness, cracks) in the pavement layers to the load repetitions (annual average daily traffic) and other affecting factors (e.g., asphalt type, weather, width of the road, layer thickness).
 Empirical models are known to suffer from limited applicability (Gupta et al.
@@
 Their analysis showed that narrower road width was noted to result in shorter pavement lifetimes, particularly in Troms, where studded tires are used one month earlier than other locations, e.g., Vestland or Trøndelag in Norway.
-Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced spatial rutting effects.
-To achieve this, we employ a hierarchical Bayesian framework comprising latent Gaussian models (LGMs) that encompass both random and spatial components (Jourdain and Klein-Paste 2024).
-Latent Gaussian models have been used in various statistical analyses, as demonstrated by their applications in a multitude of fields (e.g., Dey et al.
+1.1 Objectives
+Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced rutting.
+To achieve this, we use eleven years (2010-2020) of pavement rutting data from the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border.
+This data, collected annually, provides a robust historical record that is essential for identifying trends and anomalies in rutting patterns.
+By utilising this extensive dataset, we aim to enhance our understanding of rutting dynamics over time and across different sections of the road.
+To analyze the rutting phenomenon, we develop a hierarchical Bayesian framework utilizing latent Gaussian models (LGMs) (Jourdain et al.
+2024).
+These models are particularly suited for this type of analysis due to their flexibility and ability to explicitly account for spatial and random components (e.g., Dey et al.
 2000, Chu et al.
@@
 2009b).
-These models offer a high degree of flexibility and explicitly account for model dependence.
-They can approximate entire processes defined over continuous domains and provide probabilistic predictions (Lindgren and Rue 2015).
-In our specific case, the rutting phenomenon on road surfaces persists regardless of whether it is directly observed, with the latent fields serving as the authentic representation of this rutting.
-A crucial premise underlying spatial models is that observations situated in close proximity to one another in the spatial domain exhibit greater similarity than those positioned further apart.
-For a broad class of LGMs, the Integrated Nested Laplace Approximation approach (INLA, Rue et al.
-2009a) proves to be an apt tool for performing Bayesian inference.
-The proposed models are based on historical rutting observations and a selection of explanatory variables known to influence pavement rutting that is available for the Norwegian stretch of the road.
-These variables are pavement type, Average Annual Daily Traffic (AADT), rut depth from the previous year and lane width.
-The inclusion of these variables is grounded in existing knowledge of factors that significantly influence rutting in road surfaces (Chan et al.
+The LGMs enable us to capture the underlying processes governing rutting and provide probabilistic predictions, which are crucial for understanding and managing pavement performance.
+Advanced statistical analysis is implemented through the Integrated Nested Laplace Approximation approach (INLA) for performing Bayesian inference.
+This approach is well-suited for LGMs and will allow us to efficiently handle the complex computations involved in our models.
+By using INLA, we aim to achieve high computational efficiency and accuracy in our analysis.
+Incorporating a selection of explanatory variables known to influence pavement rutting is another key objective.
+These variables include asphalt type, Average Annual Daily Traffic (AADT), rut depth from the previous year, and lane width.
+The inclusion of these factors is based on their established impact on rutting (Chan et al.
 2010, Lang 2012, Thodesen et al.
@@
 2023).
-A critical element in our model is the role of Average Annual Daily Traffic (AADT) on the pavement.
-Rutting, as we argue, is contingent on the presence of AADT or the load imposed by traffic.
-Without traffic loading, certain forms of pavement deterioration, such as cracks and roughness, may occur, but rutting is improbable.
-This hypothesis aligns with prior research on the prediction of pavement variability and deterioration, as illustrated in the work of Hao et al.
-(2009) and Vedvik (2021).
-Rutting is a complex phenomenon, and the depth of rutting observed in the previous year is a key factor in predicting future rutting (Archilla and Madanat 2000).
-Deeper rut depths from the previous year are expected to correlate with increased rutting in the current year, and this relationship is incorporated into the model.
-Additionally, lane width is considered as a contributing factor to rutting rates.
-It is logical to assume that narrower roads are more prone to higher rutting rates due to the limited space for motorists to maneuver.
-However, it’s important to note that the effect of lane width on rutting is contingent on the intensity of traffic.
-Pavement type also plays a crucial role, particularly in colder climates where it can significantly influence rutting resistance.
-During winter, rising temperatures lead to shorter periods of frozen ground and snow cover on road surfaces, altering the frequency of freeze-thaw cycles.
-Consequently, this poses challenges for lower-grade binder types like asphalt gravel concrete (Agc), which are more susceptible to rutting under such conditions
-As proof of concept, our proposed models are applied to eleven years (2010-2020) of pavement rutting data for one road stretch; the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border.
-We identify areas with unexplained or accelerated rutting and determine which explanatory variables are most important to rutting development.
-A detailed description of the data used and data cleaning process is given in Section 2.
-Model notation, formulation and evaluation criteria are given in Section 3.
-Results and discussion are given in Section 4.
-Section 5 explains how road authorities can use the model, based on the results.
+By integrating these variables into our model, we can better account for the diverse factors that contribute to rutting, thereby improving the accuracy and reliability of our estimations.
+Identifying areas with unexplained or accelerated rutting is a crucial part of our objectives.
+By analyzing the data with our models, we pinpoint sections of the road that exhibit unexpected rutting patterns.
+Therefore, we provide practical insights for road authorities to use the model results for better pavement maintenance planning.
+By shifting from manual, suspicion-based diagnostics to a more efficient, automated approach, we hope to enable pavement managers to concentrate their diagnostic resources on identifying the correct countermeasures rather than merely locating problem areas.
+This will ultimately lead to more effective maintenance strategies and improved road safety.
 2 Materials
-2.1 Data source
-The data material used in this research is obtained from the Norwegian Road Database (NVDB) and the Rosita database.
-The NVDB is primarily developed by the Norwegian Public Roads Administration (NPRA), the road authority for national roads, including managing, researching, planning, building, and operating.
-In 2022, NVDB consisted of 93 500 kilometres of public roads, including national roads (state-owned, 10,500 kilometres), county roads (44 000 kilometres) and municipal roads (39 000 kilometres) (Stephansen 2022).
-The Rosita database, which stores extensive and calculated data from measuring systems (see Section 2.3), was made available for use in this study.
 2.1 Area of study
 […]
+Most of the road has two lanes, but within approximately 2.5 km from Stjørdal, it expands to four lanes.
+Two-lane stretches are around 6 meters wide, while four-lane sections are up to 13 meters wide.
+The expected lifetime of the road is 20-30 years after construction.
+The climate is subarctic or cold-temperate, with temperatures varying by season: winter (-2°C to 2°C), spring (5°C to 12°C), summer (15°C to 22°C), and autumn (10°C to 15°C) (Sygna and O’Brien 2001, Institutt 2010).
+On average, air surface temperatures are at most 18 °C (see Figure A.5 in Appendix A.5) .
+For analysis, we focus on one lane (lane 1) in the direction from Stjørdal to Storlien.
 2.2 Data
-The available explanatory variables for expected rut depth includes, traffic load in terms of annual average daily traffic (AADT), pavement or asphalt type (three categories), rut depth from the previous year and the lane width.
-A summary of the data used can be found in Table 1 and Figure 1 and Supplementary Material LABEL:supplementary:datasummary.
+The data for this research is sourced from the Norwegian Road Database (NVDB) and the Rosita database.
+NVDB, primarily developed by the Norwegian Public Roads Administration (NPRA), includes 93,500 kilometers of public roads, comprising national, county, and municipal roads (Stephansen 2022).
+The Rosita database provides extensive and calculated data from measuring systems (see Section 2.2.1).
+We use eleven years (2010-2020) of pavement rutting data from the European highway route E14, covering a 67.1-kilometer stretch from Stjørdal, Norway, to Storlien on the Swedish border.
+Collected annually, this data set offers a comprehensive historical record crucial for detecting trends and anomalies in rutting patterns.
+Explanatory variables for expected rutting include annual average daily traffic (AADT), asphalt type, rut depth from the previous year, and lane width.
+Table 1 summarizes the data used, and Figure 1 and Supplementary Material LABEL:supplementary:datasummary provide additional details.
 Table 1: Nomenclature and Variables
-Terms and Variables Description/Value Rut depth ($d_{i,t}$, millimetre) Measurement for year $t$ for all $20$-metre stretch ($i$); mean $=12.11$ millimetres and sd $=6.78$ millimetres Lane width (${w_{i}}$, metre) Half (or quarter) of width of 2-lane (or $2-2$ lane) used; mean $=6.02$ millimetres and sd $=0.69$ millimetres Pavement type Asphalt concrete (Ac), asphalt gravel concrete (Agc) and stone mastic asphalt (Sma) AADT (vehicles/year) $568-18500$ vehicles, with half (or quarter) of AADT for 2-lane (or $2-2$ lane) road used $r_{i,t}$ (millimetre/year) Change in rut depth between year $t$ and $t-1$ (for $t=1,...,T$).
+Terms and Variables Description/Value
+Rut depth ($d_{i,t}$, mm) Measurement for year $t$ for all $20$-metre road segments ($i$); mean $=12.11$, sd $=6.78$
+Lane width (${w_{i}}$, m) Half (or quarter) of width of 2-lane (or $2-2$ lane) used; mean $=2.99$, sd $=0.18$
+Asphalt type Asphalt concrete (Ac), asphalt gravel concrete (Agc), stone mastic asphalt (Sma)
+AADT (vehicles/day) Range: 568-18,500, adjusted for the number of lanes
+$r_{i,t}$ (mm/year) Rutting, calculated as the change in rut depth ($d_{i,t}$) between years $t$ and $t-1$
 2.2.1 Rut Depth and Rutting
-The rut depth is measured annually using a laser scanner device ViaPPS (Pavement Profile Scanner) — a measurement system developed by ViaTech in cooperation with NPRA (Peraka and Biligiri 2020).
-The annual rut depth measurement is used to calculate rutting (mm/year), the indicator variable used for assessing pavement performance.
+Rut depth is measured annually using the ViaPPS (Pavement Profile Scanner - laser scanner), developed by ViaTech in collaboration with the NPRA (Peraka and Biligiri 2020).
+Measurement positions are determined with high-accuracy GPS (CPOS/DPOS — satellite-based position service with centimetre/decimetre accuracy), ensuring precise location data.
+The rut depth is estimated by comparing the height of the ridges formed between the wheel paths, which involves identifying the two lowest points in the wheel path and measuring the distance to the top of the ridge.
+Additionally, the laser covers an approximate 4-meter footprint, ensuring that the vehicle remains within the wheel tracks and avoids crossing the centerline or shoulder.
+Thus, minor deviations (10-50 cm) do not affect the calculations.
+Annual rut depths along the road range from 0 to 41 mm, measured at 20-metre intervals.
+There are 3355 road segments measured for annual rut depth.
+Rutting, denoted as $r_{i,t}$ (mm/year), is calculated as the difference in rut depth between years, as shown in Equation 1.
 We denote the rut depth variable at road segment $i=1,...,n$ in year $t$ as $d_{i,t}$, and $r_{i,t}$ as rutting (mm/year).
-Rutting for year $t$ at segment $i$ is defined as
+Rutting for year $t=1,...,T$ at segment $i$ is defined as
 $r_{i,t}=d_{i,t}-d_{i,t-1}.$ (1)
-In the case of repavement, rut depth is smaller in the proceeding year resulting in large negative values of $r_{i,t}$. Therefore, following Vedvik (2021), large negative rutting values are removed;
-$r_{i,t}=\begin{cases}r_{i,t},&\text{if }r_{i,t}\geq-d_{i,t}/2\\ \text{NaN }&\text{otherwise, }\end{cases}$ (2)
-i.e., $r_{i,t}$ is kept if it is at least half the rut depth, or assigned NaN (not a number) otherwise.
-Some missing rutting data may occur due to uncertainty arising from the actual rut depth measurements or the data collection process (Figure 1, graph (c) shows large variability in rutting).
+When repavement occurs, rut depth decreases, sometimes resulting in large negative values of the rutting ($r_{i,t}$).
+Following Vedvik (2021), large negative rutting values are considered missing and set to NaN (not a number):
+$r_{i,t}=\begin{cases}r_{i,t}&\text{if }r_{i,t}\geq-\frac{d_{i,t}}{2}\\ \text{NaN}&\text{otherwise,}\end{cases}$
+and the small negative values kept in the dataset are considered as measurement uncertainties and are accounted for in the random variability.
+Some missing rutting data may occur due to measurement uncertainty or data collection process.
+For available observations, the overall mean rutting and segment mean rutting (over all years) are calculated as $\bar{r}={1}/{n}\sum_{\forall t}\sum_{\forall i}r_{i,t}$ and $\bar{r}_{i}={1}/{T}\sum_{t=1}^{T}r_{i,t}$, respectively.
+2.2.2 Lane Width
+In the database, road width and the number of lanes are available.
+The first 2.5 km from Stjørdal has four (4) lanes, the remaining stretch has two lanes.
+Lane width is calculated by dividing the road by the number of lanes.
 2.2.3 Annual Average Daily Traffic (AADT)
-The annual average number of daily traffic (AADT) f is available for each road segment for each year.
-AADT ranges from 568-18500 between the years $2010-2020$, with more traffic close to Stjørdal and less towards Sweden.
-NPRA states that high-traffic roads have an AADT of more than 5000 vehicles and recommends maintenance at rut depth values exceeding $20$ millimetres, while low-traffic roads with AADT lower than 5000 require pavement treatment when rut depth exceeds $25$ millimetres.
-Up to $2.5$ kilometres along the road, a quarter of the AADT is used in the analysis or half of the AADT otherwise.
-In all analyses, AADT in ten thousands are used.
-2.3.3 Pavement type
-In Norway, asphalt concrete is used on all types of roads and in all traffic classes.
-The asphalt mixtures used on the European route E14 are; asphalt concrete mixtures (Ac) and asphalt gravel concrete (Agc), and a stone mastic asphalt mixture (Sma).
-These include stone mastic asphalt Sma is the stronger and densely graded asphalt mass with unmodified binders and is mainly used on high-traffic roads; and asphalt concrete (Ac) and asphalt gravel concrete (Agc), which are mainly used on low-traffic roads of up to 5000 vehicles, courtyards, sidewalks and bicycle paths.
-These are softer, dynamic pavement mixtures, with less stringent requirements for the nature and grading of the stone material compared with that of stone mastic asphalt.
-Asphalt concrete, is however, more stringent that Asphalt gravel concrete; it is a dense-graded asphalt used as a protective layer over the road surface, or laid over a layer of asphalt gravel concrete (d’Angelo et al.
-2008, vegvesen Vegdirektoratet 2009, Taddesse 2013, Zhang et al.
-2022).
-2.3.4 Lane width
-The road width is quite variable and depends on the number of lanes.
-Since one lane is considered for the purposes of analysis, a quarter of the measured width up to 2.5 kilometres from Stjørdal or half of the measured width for the rest of the road stretch is used.
-(a) AADT/Pavement type, and lane width
+The Annual Average Daily Traffic (AADT) for each road segment between 2010 and 2020 varied from 568 to 18,500 vehicles per day (Figure 1, graph (a)).
+Fluctuations in traffic volume over short distances are influenced by various factors.
+Notably, there are considerably larger traffic volumes within Stjørdal to approximately 2.5 km towards Storlien.
+The European route E14 runs through the city center of Stjørdal, contributing to high traffic levels in the initial few kilometers.
+The Norwegian Public Roads Administration (NPRA) recommends maintenance based on AADT values.
+High-traffic roads should be treated when rut depths exceed 20 mm, while low-traffic roads should be treated when rut depths exceed 25 mm.
+Adjusting AADT values based on road width is essential.
+Throughout the analyses, AADT values are expressed in ten-thousands.
+2.2.4 Asphalt Type
+Asphalt mixtures in cold-temperature regions like Norway are specifically designed to withstand harsh climatic conditions.
+These mixtures may include Polymer Modified Asphalt (PMA) to prevent cracking due to freeze-thaw cycles and low temperatures, and low-temperature performance grade binders to maintain flexibility at sub-zero temperatures.
+Dense-graded aggregates are used to minimize water infiltration, crucial for preventing freeze-thaw damage.
+The primary focus is on enhancing flexibility, low-temperature performance, and resistance to moisture, employing softer binders, dense-graded aggregates, and anti-stripping agents (Snilsberg et al.
+2006, d’Angelo et al.
+2008, Saba et al.
+2009, Ødegård 2015).
+Asphalt concrete (Ac), asphalt gravel concrete (Agc), and stone mastic asphalt (Sma) are the primary surface mixtures used in Norway.
+Sma, known for its superior wear resistance and deformation resistance due to the larger- and high-quality-stone in its composition, is typically employed on high-traffic roads.
+In contrast, Ac and Agc are commonly used on low-traffic roads, courtyards, sidewalks, and bicycle paths.
+(a) AADT/Asphalt type and lane width
 (b) Rutting
-Figure 1: AADT together with pavement type, and lane width (a), and rutting (for years 2019-2020, mm/year — missing in red) (b).
+Figure 1: AADT together with asphalt type, and lane width (a), and rutting for a short stretch of E14 (for years 2019-2020, mm/year — missing in red) (b).
 2.3 Exploratory analysis of rutting
-On average, over all stretches and years, rutting is $1.62$ mm/year; see Figure  2 (a).
-The density plots,  2 graph (b), show rutting for all segments for each year.
+On average, over all stretches and years, the overall rutting is $1.62$ mm/year; see Figure  2 (a).
+The density plots, Figure  2 graph (b), show rutting for all segments for four selected years (2017-2020).
 Some years having higher rutting levels and are more widely distributed than others.
-The lines in graphs (c) and (d) give the correlation of rutting values with respect to distance (mean across all years, and two randomly selected years; 2011 and 2012, respectively) (Bachmaier and Backes 2011).
-It is computed from the observed rutting and shows how the rutting values are correlated with respect to distance between the corresponding segments.
-The distance (metre) is shown on the x-axis, and the calculated correlation between rutting values is on the y-axis.
-The graph shows some spatial dependencies of pavement rutting along the road.
-These exist for short distances of up to ($\sim$) $300$ metres (3 kilometres).
-Refer to caption (a) Observed rutting Refer to caption (b) Density Refer to caption (c) Autocorrelation across years Refer to caption (d) Autocorrelation for years 2011-2012
-Figure 2: Graph (a): Observed rutting for each year (2011-2020), observed mean rutting $1.62$ mm/year (     ), and the observed mean rutting across segments ( ).
-Graph (b): Rutting density plots, graph.
+The lines in graphs (c) and (d) represent the empirical correlation of rutting values with respect to distance.
+Specifically, graph (c) shows the segment mean rutting over all years (see Section 2.2.1), and graph (d) displays the correlation for two selected years (2011-2012).
+These values are computed from the observed rutting at 20-meter segments, illustrating the empirical correlation as a function of the distance between the segments (Székely et al.
+2007, Jentsch et al.
+2020).
+Based on the data, we present results for distances up to 300 metres.
+In graph (d), the x-axis represents the distance in meters, while the y-axis shows the calculated correlation between rutting values.
+This graph highlights the spatial dependencies of pavement rutting along the road for specific years.
+When compared to graph (c), which reveals long-term spatial patterns and trends, graph (d) demonstrates that the correlation of rutting values tends to diminish more quickly over shorter distances.
+Graph (c) indicates that there are meaningful correlations in rutting values for distances up to approximately 150 meters, suggesting that these measurements capture underlying trends that persist over longer stretches of the road.
+In contrast, graph (d) shows that for specific years, the correlation between rutting values is important only up to about 100 meters.
+This shorter correlation distance implies that the rutting patterns observed in a given year are influenced by more localized factors and exhibit greater variability over shorter distances.
+The difference in correlation distances between the two graphs shows the importance of considering both short-term and long-term spatial dependencies when analyzing pavement conditions.
+While long-term data provide a comprehensive view of how road conditions evolve, short-term data, as depicted in graph (d), highlight the influence of yearly variations and localized factors on pavement rutting.
+This dual approach enables a better understanding of the factors affecting road quality and can inform more effective maintenance and rehabilitation strategies.
+(a) Observed rutting
+s Refer to caption
+(b) Density plot of rutting
+(c) Autocorrelation of segment mean rutting
+(d) Autocorrelation for years 2011-2012
+Figure 2: Graph (a): Observed rutting for each year (2017-2020), the overall mean rutting ($\bar{r}=1.62$ mm/year (for all years, 2011-2020) (     ), and segment mean rutting ($\bar{r}_{i}$, over all 11 years) ( ).
+Graph (b): Density plots for rutting.
 Graphs (c) and (d) show how rutting values are correlated with distance.
-Spatial dependencies in rutting exist up to 3 kilometres.
+Spatial dependencies in rutting exist up to 150 metres.
 3 Latent Gaussian models for Rutting
 A Bayesian latent Gaussian model is a three-stage hierarchical model consisting of a likelihood, a latent Gaussian field, and hyperparameters.
-In this paper, the true rutting at location $s_{i}$ in year $t$ is denoted by $\eta_{t}(s_{i})$ and it is modelled as a latent Gaussian field that includes explanatory variables and spatial effects.
+In this paper, the modelled rutting (i.e., all rutting behavior) at location $s_{i}$ in year $t$ is denoted by $\eta_{t}(s_{i})$ and it is modelled as a latent Gaussian field that includes explanatory variables and spatial effects.
 We specify six (6) models with different explanatory variables included and with and without spatial and random effects.
-3.1 Likelihood: observed rutting given true rutting
-We denote the observed rutting for segment $i$ in year $t$ as $r_{i,t}$. The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as
+3.1 Likelihood: observed rutting given modelled rutting
+We denote the observed rutting (detected and measured) for segment $i$ in year $t$ as $r_{i,t}$.
+The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as
 $r_{i,t}\ |\ \eta_{t}(s_{i})\sim\mathrm{N}\left(\eta_{t}(s_{i}),\sigma^{2}_{\varepsilon}\right).$ (2)
@@
 The unstructured error terms $\varepsilon_{i,t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\mathrm{Normal}(0,\sigma^{2}_{\varepsilon})$ are independent and identically distributed (i.i.d).
-3.2 Latent field: model for true rutting
+3.2 Latent field: modelled rutting
 We model the observed rutting as Gaussian with expected value $\eta_{t}(s_{i})$ and variance $\sigma^{2}_{\varepsilon}$.
@@
 It is known that more traffic gives more rutting, so we include traffic (AADT) in the model as an interaction with each of the explanatory variables.
-This means that for pavement type, we assume that the rutting is linear for all pavement types (double traffic means double rutting), but that the different pavement types can have different rates of deterioration.
+This means that for asphalt type, we assume that the rutting is linear for all asphalt types (double traffic means double rutting), but that the different asphalt types can have different rates of deterioration.
 From the explanatory analysis (Figure 2), we found that the rutting seems to be on different levels for different years, and this might be explained by different weather conditions.
@@
 $\centering\begin{split}\eta_{t}(s)&=\beta_{\mathrm{Ac}}z_{\mathrm{Ac}(s,t)}+\beta_{\mathrm{Agc}}z_{\mathrm{Agc}(s,t)}+\beta_{\mathrm{Sma}}z_{\mathrm{Sma}(s,t)}+\beta_{\mathrm{d-1}}z_{\mathrm{d-1}(s,t)}+\beta_{\mathrm{w}}z_{\mathrm{w}(s,t)}+\gamma_{t}.\end{split}\@add@centering$ (3)
-The model assumes that the level of expected rutting (mm/year) is influenced by an interaction effect of the annual average daily traffic (AADT) and some explanatory variables, including the level of rut depth from the previous year ($z_{\mathrm{d-1}}$), the lane width ($z_{\mathrm{w}}$) and pavement type.
-Technically, $z_{Ac}(s,t)$ is
-$\begin{cases}\text{AADT (in ten thousands) if pavement at ($s,t$)is Ac,}\\ 0\ \text{if not Ac;}\\ \end{cases}$
-similarly, for pavement types Agc and Sma, representing asphalt gravel concrete and stone mastic asphalt, respectively.
-The effect of lane width, say, on rutting is dependent on the traffic intensity.
+The model assumes that the level of expected rutting (mm/year) is influenced by an interaction effect of the annual average daily traffic (AADT) and some explanatory variables, including the level of rut depth from the previous year ($z_{\mathrm{d-1}}$), the lane width ($z_{\mathrm{w}}$) and asphalt type.
+This assumption stems from our prior understanding: when there’s no traffic, there’s no rutting (or rut depth).
+This insight is grounded in research, such as the work of Hao et al.
+(2009), which identified increased loading time (traffic loads) as the primary factor accelerating rutting in asphalt pavements.
+Let $z_{\mathrm{AADT}(s,t)}$ denote AADT for location $s$ in year $t$, and $I_{\mathrm{Ac}(s,t)}$ be an indicator variable that has a value of 1 if location $s$ in year $t$ has asphalt type Ac and 00 otherwise.
+Then, $z_{\mathrm{Ac}(s,t)}=I_{\mathrm{Ac}(s,t)}\cdot z_{\mathrm{AADT}(s,t)}$ and, hence, equals AADT if the asphalt type is Ac (asphalt concrete) and 00 otherwise.
+Similarly, $z_{\mathrm{Agc}(s,t)}$ and $z_{\mathrm{Sma}(s,t)}$ equal the AADT for location $s$ in year $t$ if the asphalt type is Agc (asphalt gravel concrete) and Sma (stone mastic asphalt), respectively, and 00 otherwise.
+Further, $z_{\mathrm{d-1}(s,t)}=d(s,t)\cdot z_{\mathrm{AADT}(s,t)}$ and $z_{\mathrm{w(s,t)}}=\mathrm{w}{(s,t)}\cdot z_{\mathrm{AADT}(s,t)}$.
 The random yearly effect is assumed to be distributed as $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},...,\gamma_{T}){\sim}\mathrm{Normal}(\boldsymbol{0}_{T},\ \tau_{\gamma}^{-1})$.
 3.2.2 Models including spatial effects
-There are several variables not included in our model (Equation (4)) that are known to influence rutting.
-Examples are variables related to weather, bearing capacity and proportion of heavy traffic load.
-Most of these variables are spatially dependent, and segments close in space are likely to be more similar than segments further a part.
-This can be accounted for and utilised by including spatial random effects.
-The spatial random effect is a Gaussian process (GP) along the road and can be interpreted as the part of the model not explained by the explanatory variables; lane width, rut depth from the previous year and pavement type that is spatially dependent.
-Further, it is reasonable that some of the spatial random effect is consistent over years (e.g., bearing capacity, side slope height, surface and base curvature and proportion of heavy traffic) while others vary between years (e.g., heavy rain events or more heavy traffic due to construction work).
-Therefore, we include both an annually varying spatial process $\xi_{t}(\boldsymbol{s})$ for year $t=1,...,T$, and a common spatial process $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component.
-The spatial rutting component is the spatial contribution to the total rutting that is not explained by the other variables e.g., variables in the non-spatial model in Equation (4).
-This contribution can either be positive or negative.
+Our model (Equation (3)) does not include several variables known to influence rutting, such as weather, bearing capacity, and proportion of heavy traffic load.
+These variables are spatially dependent, with segments closer in space likely to be more similar than those further apart.
+To account for this spatial dependency, we incorporate a Gaussian process (GP) as a spatial random effect along the road.
+GPs are powerful tools for modeling spatial dependencies, capturing variability due to both known factors (e.g., road characteristics, traffic patterns, and other explanatory variables) and unknown factors not explicitly accounted for in the model.
+The unknown factors could include unmeasured road characteristics, environmental variables like temperature, or other spatially dependent factors influencing rutting.
+Our model aims to capture both consistent spatial effects that remain stable over time and spatial effects that vary between years.
+Consistent spatial effects, such as road characteristics (e.g., bearing capacity, side slope height, surface and base curvature, and proportion of heavy traffic), are likely to influence rutting consistently over time.
+Additionally, there are spatial effects that vary between years, such as heavy rain events, temperature variations, or fluctuations in heavy traffic due to construction work.
+To account for these variations, we include both an annually varying spatial process, $\xi_{t}(\boldsymbol{s})$, for each year $t=1,...,T$, and a common spatial process, $\omega(\boldsymbol{s})$, referred to as the spatial rutting component.
+The annually varying spatial process, $\xi_{t}(\boldsymbol{s})$, captures the spatial variability in rutting that changes from year to year, accounting for spatial variations specific to each year, such as the impact of specific weather events or construction activities.
+The common spatial process, $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component, captures the spatial variability in rutting that remains consistent across all years.
+This includes spatial effects not explicitly accounted for by other variables in the model, such as road characteristics and other spatially dependent factors.
+This contribution can be either positive or negative.
 The linear predictor for the model including spatial effects is described as
@@
 where the remaining model parameters are defined as that in Equation 3.
-The common spatial process $\omega(\boldsymbol{s})$ in Equations (5) is the time constant variable that can give rutting not explained by known variables (or unexplained rutting).
+The common spatial process $\omega(\boldsymbol{s})$ in Equations (4) is the time constant variable.
 It is a Gaussian random process $\left\{\omega(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{D}\subseteq\mathbb{R}^{d}\right\}$ such that for any $n\geq 1$ and for each set of spatial locations $\left(s_{1},...,s_{n}\right)$ satisfies
@@
 4 Results and discussion
-We fit the models in Section 3.2.3 to the data in Section 2.3.
+We fit the models in Section 3.2.3 to the data in Section LABEL:sec:variables.
 For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation.
@@
 4.1 Estimated spatial rutting component and yearly effect
-We analyze estimates from Model 1, identifying areas with high rutting rates not explained by AADT, pavement type, and lane width.
+We analyze estimates from Model 1, identifying areas with high rutting rates not explained by AADT, asphalt type, and lane width.
 The spatial rutting component ()$\hat{\omega}(s)$) plotted in Figure 3(a), varies significantly along the road.
@@
 Rutting varies notably between years for spatial and non-spatial models, with minimum estimates in 2016 and maximum in 2018.
-A systematic shift between spatial and non-spatial models is evident, particularly in pavement type coefficients (see Table 3 in Section 4.2).
+A systematic shift between spatial and non-spatial models is evident, particularly in asphalt type coefficients (see Table 3 in Section 4.2).
 Including spatial components in the model enhances rutting variability explanation and identifies areas with unexplained high rutting rates.
@@
 The estimated parameters and 95% credible intervals (CI) for spatial (Model 1) and non-spatial (Model 4) are shown in Table 3.
-Spatial models exhibit notably higher estimates for pavement type coefficients compared with non-spatial models (see also Table A.2 in Appendix A).
+Spatial models exhibit notably higher estimates for asphalt type coefficients compared with non-spatial models (see also Table A.2 in Appendix A).
 Moreover, models lacking the spatial rutting component display larger variance in the ‘noise’ term ($\hat{\sigma}{\varepsilon}$), suggesting that the spatial rutting component captures unobservable spatially varying factors such as weather and traffic.
@@
 4.2.1 Effects of the explanatory variables
-Pavement types, particularly those with weaker binder concentrations like asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) and asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$), experience more rutting (Table 3).
+Asphalt types, particularly those with weaker binder concentrations like asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) and asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$), experience more rutting (Table 3).
 Qian et al.
 (2019) and Mascarenhas et al.
-(2020) shed light on how temperature and heavy traffic ($>$ 3.5 tonnes) impact rutting resistance on pavement types such as Agc and Ac.
+(2020) shed light on how temperature and heavy traffic ($>$ 3.5 tonnes) impact rutting resistance on asphalt types such as Agc and Ac.
 Notably, asphalt gravel concrete is predominantly situated approximately 30 kilometers from the Swedish border, where a sub-arctic climate, underdeveloped networks, and heavy traffic converge to adversely affect pavement performance Peel et al.
@@
 There is also greater uncertainty in asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), although considering spatial dependencies helps to explain more of the variability.
-Non-spatial models appear to explain variability between years rather than by pavement type, suggesting a nuanced relationship between these factors in understanding rutting susceptibility.
-Rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes more crucial without accounting for spatial variability (Model 4, Table 3).
+Non-spatial models appear to explain variability between years rather than by asphalt type, suggesting a nuanced relationship between these factors in understanding rutting susceptibility.
+The rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes increasingly important when spatial variability is not considered (Model 4, Table 3).
+As a genuine spatial-temporal variable, including the previous year’s rut depth in the model modifies the spatial effect $\hat{\sigma}_{\omega}(s)$, specifically reducing it, which in turn enhances the precision of the estimates by decreasing uncertainty, $\hat{\sigma}_{\varepsilon}$.
+This improvement is evident in Models 1 and 3 (see Table 4 in Section A).
+Incorporating additional relevant spatial-temporal variables could further refine the results, providing a better explanation for the observed uncertainty.
 The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is minimal and positive, suggesting negligible impact on expected rutting.
-These findings indicate that changes in coefficients are significant when spatial dependencies of unmodeled factors (e.g., weather) are considered.
-Furthermore, the spatial model captures more uncertainty (smaller $\hat{\sigma}_{\varepsilon}$).
 4.3 Estimated rutting from the explanatory variables, random yearly effect and spatial rutting component
@@
 (b) $\hat{\eta}$, AADT and lane width
-Refer to caption (c) $\hat{\eta}>2$ mm Refer to caption (d) $\hat{\eta}>3$ mm.
-Figure 4: Graph (a): Expected rutting from: i) $\hat{\eta}$ and ii) from $*\hat{\eta}{{}^{*}}$, with each asphalt type highlighted.
+(c) $\hat{\eta}>2$ mm
+(d) $\hat{\eta}>3$ mm.
+Figure 4: Graph (a): Expected rutting from: i) $\hat{\eta}$ and ii) from $\hat{\eta}{{}^{*}}$, with each asphalt type highlighted.
 The effects are inferred for the year 2020.
 Graph (b): Expected rutting plotted against AADT and lane width for a short stretch.
-Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year and $3$ mm/year, respectively.
+Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year and $3$ mm/year, respectively
 4.3.1 Comparing expected rutting with observed data
@@
 Additionally, the observed mean rutting across segments ranges from -0.14 to 8.33 mm/year (Figure 2, graph (a)), while the maximum expected rutting from the model is 4.67 mm/year (Figure 4, graph (a)), approximately half the maximum observed mean rutting.
-In Figure 4, graph (b), we plot the expected rutting $\hat{\eta}$, lane width, and AADT (per ten thousand) for a short road section spanning segments $1500-2700$ (or $30-54$ km from Stjørdal), characterized by weaker pavement type Agc.
+In Figure 4, graph (b), we plot the expected rutting $\hat{\eta}$, lane width, and AADT (per ten thousand) for a short road section spanning road stretch 42000 to 54000 metres (or $42-54$ km from Stjørdal), characterized by weaker asphalt type Agc.
 This section exhibits spikes of accelerated rutting, particularly evident with increasing traffic.
@@
 The significant spike in Figure (4) (a), indicating high rutting rates, prompts further investigation.
-We examine road segments spanning 2100 to 2700 meters surrounding this spike, located in the municipality of Meråker, just 20 kilometers west of the Swedish border at Storlien.
+We examine the road stretch spanning 42000 to 54000 metres surrounding this spike, located in the municipality of Meråker, just 20 kilometers west of the Swedish border at Storlien.
 This stretch, depicted in Figure 4 (a), exhibits notably high rutting, peaking at $4.67$ mm, with an estimated lifespan of only $5$ years (in a low-traffic area requiring maintenance at 25 mm rut depth).
-Possible contributors to this accelerated rutting include clogged drainage, flooding, absence of side ditches (see Figure 5), heavy traffic, or road designs featuring railguards installed close to the lane edge, which encourages traffic to remain within the same wheel tracks (refer to Figure A.4 in Appendix A.4).
+Possible contributors to this accelerated rutting include clogged drainage, flooding, absence of side ditches, heavy traffic, or road designs featuring railguards installed close to the lane edge, which encourages traffic to remain within the same wheel tracks (refer to Figure A.4 in Appendix A.4).
 Meråker’s industrial and agricultural activities result in increased traffic from heavy vehicles, including farm equipment such as tractors (industrial, wheel drive, row crop).
-(a) Short section between EV14 S6D1 m3841-m4285 F1 with high rutting.
-(b) Location EV14 S6D1 m4068 F1 in year 2020
-(c) Location EV14 S6D1 m4059 F1 in year 2021
-(d) Location EV14 S6D1 m4065 F1 in year 2022
-Figure 5: Graph (a): Location with high rutting rates (road segments $2100-2700$ towards Sweden).
-Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@335074,7038866,15/vegsystemreferanse:335035.067:7038844.352
 4.4.1 Exploring the spatial model with various values of the explanatory variables
@@
 Second, we consider alternative asphalt types - asphalt concrete (Ac) or stone mastic asphalt (Sma) ($\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$).
-Third, we examine extreme values such as the strongest pavement (Sma), minimal rut depth (0.01 mm), or wider lanes (6 meters) ($\hat{\eta}_{\mathrm{Best}}$).
+Third, we examine extreme values such as the more wear- and deformation-resistant pavement (Sma), minimal rut depth (0.01 mm), or wider lanes (6 meters) ($\hat{\eta}_{\mathrm{Best}}$).
 Conversely, accelerated rutting scenarios involve deeper rut depth (25 mm), narrower lanes (2.75 meters), and asphalt gravel concrete (Agc) ($\hat{\eta}_{\mathrm{Worse}}$) from the E14 dataset.
@@
 Using the best covariate values ( ), rutting is projected to be at most 4.30 mm/year, with maintenance required approximately 6 years after construction or maintenance.
-Additionally, using the stronger asphalt binder Sma yields similar rutting estimates but generally lower than those using the best values, indicating asphalt type as a primary factor.
+Additionally, using the asphalt binder Sma yields similar rutting estimates but generally lower than those using the best values, indicating asphalt type as a primary factor.
 Asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) show similar rutting estimates due to similar binder concentrations.
-(a) Expected rutting at road segments $2100-2500$ (metres) towards Sweden.
+(a) Expected rutting at road segments $42000-54000$ (metres) towards Sweden.
 (b) Expected rutting from adding each explanatory variable added sequentially.
-(c) Bearing capacity (tonn).
+(c) Bearing capacity (ton).
 (d) SCI/BCI.
 (e) SCI/BCI for the whole E14.
-Figure 6: Graph (a) displays expected rutting using spatial rutting component $\hat{\omega}(s)$ for road segments $2100-2700$, along with the ’best’ or ’worst’ explanatory values.
-$\hat{\eta}_{\mathrm{Agc}}$ represents pavement type Agc, random yearly effect $\hat{\gamma}$, and $\hat{\omega}(s)$.
+Figure 5: Graph (a) displays expected rutting using spatial rutting component $\hat{\omega}(s)$ for road stretch 42000 to 54000 metres, along with the ‘best’ or ‘worst’ explanatory values.
+$\hat{\eta}_{\mathrm{Agc}}$ represents asphalt type Agc, random yearly effect $\hat{\gamma}$, and $\hat{\omega}(s)$.
 Graph (b) illustrates expected rutting inferred from $\hat{\omega}(s)$, $\hat{\gamma}$, and sequentially added explanatory variables.
 Graph (c) shows 2021 bearing capacity in tons; values $\geq$ 16 indicate good parameters.
-Graphs (d) and (e) depict SCI to BCI ratios for a short stretch or the entire E14; values $<$ 12 indicate good curvature index.
-We examine the impact of pavement type on rutting (Figure 6, graph (b)).
-Changing from asphalt type Agc (asphalt gravel concrete) to the stronger Sma (stone mastic asphalt) while keeping all other variables constant, we sequentially add additional explanatory variables to the model, starting with pavement type.
-Initially, pavement type (Agc or Sma) is included with the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$ as follows: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$.
+Graphs (d) and (e) depict SCI to BCI ratios for a short stretch or the entire E14, where values $<$ 12 indicate good curvature index.
+Maximum y-value plotted is 50 (values $>50$ are denoted by the red ‘x’ symbol).
+We examine the impact of asphalt type on rutting (Figure 5, graph (b)).
+Changing from asphalt type Agc (asphalt gravel concrete) to the more wear- and deformation-resistant Sma (stone mastic asphalt) while keeping all other variables constant, we sequentially add additional explanatory variables to the model, starting with asphalt type.
+Initially, asphalt type (Agc or Sma) is included with the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$ as follows: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$.
 Subsequently, the previous year’s rut depth is incorporated, represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$.
@@
 Results from using stone mastic asphalt (Sma) are depicted by solid lines, and asphalt gravel concrete (Agc) results are shown with dotted lines.
-The findings affirm that pavement type significantly influences rutting.
-Implementing the stronger binder type Sma instead of the actual pavement type Agc would decrease expected rutting, reaching a maximum of $4.35$ mm/year (compared to 4.68 mm/year), approximately reducing rutting by 0.33 mm/year annually.
+The findings affirm that asphalt type significantly influences rutting.
+Implementing the binder type Sma instead of the actual asphalt type Agc would decrease expected rutting, reaching a maximum of $4.35$ mm/year (compared to 4.68 mm/year), approximately reducing rutting by 0.33 mm/year annually.
 Over a projected lifetime of 6 years, this equates to an additional year of serviceability.
 4.4.2 Evaluating pavement subgrade support on expected rutting
-We examine bearing capacity, surface curvature index (SCI), and base curvature index (BCI) for road segments 2100-2700 using 2021 data, indicative of construction quality and soil pressure sustainability over the past decade (Figure 6, graph (c)).
-Bearing capacity values exceeding 16 tons (       ) suggest a robust foundation design.
+We examine bearing capacity, surface curvature index (SCI, micrometre), and base curvature index (BCI) for road segments 42000-54000 using 2021 data, indicative of construction quality and soil pressure sustainability over the past decade.
+Bearing capacity is a measure of the deflection caused by a rolling load axle on pavement.
+It is measured using the Rapid Pavement Tester (Raptor, https://www.ramboll.com/rst/rst-products), which continuously collects data on the pavement structure.
+A value exceeding 16 metric tons (       ) suggests a robust foundation design (Figure 5, graph (c)).
 The SCI to BCI ratio (Figure 5, graphs (d) and (e)) reveals deflection location within the pavement layers.
@@
 Bearing capacity data, sampled and temperature-corrected during summer, may inflate measurements compared to thawing periods.
-Varying construction material quality along the road stretch (segments 1500-3355) warrants investigation into surface and subsoil layer deflections.
+Varying construction material quality along the road stretch (30000-67100 metres) warrants investigation into surface and subsoil layer deflections.
 An SCI/BCI ratio below 12 (       ) indicates robust construction.
 Inconsistent ratios suggest layer deformation, with data indicating uneven construction on E14, featuring deeper base layer deflections.
-In conclusion, pavement type significantly influences rutting, recommending stronger binder types for extended lifespan.
+In conclusion, asphalt type significantly influences rutting, recommending more wear resistant and deformation resistant binder types for extended lifespan.
 5 How road authorities can make use of this model
@@
 Additionally, the model provides a comprehensive understanding of rutting, enhancing authorities’ ability to make informed decisions about maintenance priorities and pavement materials during construction or rehabilitation.
-The model can investigate contributing factors to rutting beyond pavement type, such as drainage issues, heavy traffic, and road design flaws.
+The model can investigate contributing factors to rutting beyond asphalt type, such as drainage issues, heavy traffic, and road design flaws.
 Comparisons with non-spatial models emphasize the importance of spatial factors in explaining rutting variations, reinforcing the need for spatial models for accurate inference and predictions.
@@
 6 Conclusion
-This research offers valuable insights into the rutting phenomenon by elucidating both well-established factors and those that remain unexplained.
-It shows the importance of incorporating spatial information to accurately estimate rutting variability, showing that observations in close proximity exhibit higher similarity than those at a distance.
-To achieve a comprehensive understanding, it is imperative to consider both observable elements like traffic and temperature, as well as unobservable factors like weather, and integrate spatial rutting components.
-Moreover, the study confirms the significant impact of asphalt type selection on rutting performance, with rut-resistant mixtures proving more effective in mitigating rutting.
-Lane width also influences pavement rutting, with narrower roads experiencing more concentrated traffic loads and, consequently, increased rutting.
-Conversely, wider roads distribute traffic loads over a larger area, reducing rut formation.
-The analysis reveals that wider road sections tend to experience higher rut depths compared to narrower road stretches.
-This result can be attributed to the increased traffic volumes and higher wheel loads present on wider roads, which contribute to more pronounced rutting.
-To address this issue and minimize rutting problems over time, it is advisable to integrate a life-cycle cost-benefit perspective into road design and planning processes.
-Further research is encouraged to explore additional factors influencing rutting, such as bearing capacity, ditch depth for water outflow, and lane width.
-This would enhance the credibility of research results and the applicability of models to other road networks.
-Additionally, evaluating the long-term performance of rut-resistant asphalt mixtures under various traffic and environmental conditions would be worthwhile.
+This research provides valuable insights into the rutting phenomenon by elucidating both well-established factors and those that remain unexplained.
+It demonstrates the importance of incorporating spatial information to accurately estimate rutting variability, showing that observations in close proximity exhibit higher similarity than those at a distance.
+To achieve a comprehensive understanding, it is imperative to consider both observed quantities as well as unobserved quantities.
+Due to Norway’s specific climatic context, where temperatures rarely exceed 30 ${}^{\circ}\mathrm{C}$ and annual averages remain below 20 ${}^{\circ}\mathrm{C}$, temperature was not included in our rutting model.
+The maximum average mean temperature per month is less than 18 ${}^{\circ}\mathrm{C}$, indicating that temperature extremes, which could significantly impact rutting by accelerating asphalt softening, are infrequent.
+Therefore, we focused on more relevant factors such as asphalt type, prior rut depth, and traffic load, which exhibit significant spatial and temporal variability.
+This approach provides a more accurate and contextually appropriate understanding of the rutting phenomenon in Norway.
+The study confirms the significant impact of asphalt type selection on rutting performance, with rut-resistant mixtures proving more effective in mitigating rutting.
+It also highlights the influence of lane width on pavement rutting.
+Narrower roads experience more concentrated traffic loads, leading to increased rutting compared to a wider road that is otherwise similar (i.e., same AADT, asphalt type and rut depth levels).
+However, wider roads, while distributing traffic loads over a larger area, tend to experience higher rut depths, likely due to increased traffic volumes and higher wheel loads.
+To address rutting and minimize its effects over time, integrating a life-cycle cost-benefit perspective into road design and planning processes is advisable.
+Further research should explore additional factors influencing rutting, such as bearing capacity, ditch depth for water outflow, and lane width, to enhance the credibility of research results and model applicability to other road networks.
+Additionally, evaluating the long-term performance of rut-resistant asphalt mixtures under various traffic and environmental conditions would be beneficial.
+This comprehensive approach ensures that our model provides more meaningful insights and practical recommendations for mitigating rutting in similar climatic regions.
 Acknowledgement
@@
 Table 4 compares estimates from the proposed spatial and non-spatial models.
-The results show that the spatial models account for uncertainty in factors that vary in space, e.g., weather and other explanatory variables such as bearing capacity and roadside ditches not included in the model.
-Including more explanatory variables reduces the ‘noise‘ term ($\hat{\sigma}_{\varepsilon}$) and random yearly effect ($\hat{\sigma}_{\gamma}$).
+The spatial models account for spatially varying uncertainties, such as weather and explanatory variables, e.g., bearing capacity.
+Including more explanatory variables reduces the noise term ($\hat{\sigma}_{\varepsilon}$) and the random yearly effect ($\hat{\sigma}_{\gamma}$).
 Table 4: Estimated model parameters with $95\%$ credible intervals
@@
 This is because asphalt type is spatially varying.
-Stone mastic asphalt (Sma) gave lower rutting values because it has a stronger binder.
-(a) Pavement type
+Stone mastic asphalt (Sma) gave lower rutting values because it has a more wear resistant binder.
+(a) Asphalt type
 (b) Rut depth
-Figure A.2: Marginal posterior distributions of pavement type (solid lines are the spatial models, and the dotted lines are the non-spatial models), rut depth from the previous year and lane width.
+Figure A.2: Marginal posterior distributions of asphalt type (solid lines are the spatial models, and the dotted lines are the non-spatial models), rut depth from the previous year and lane width.
 A.3 Marginal Posterior Distributions of Random Effects
-Figure A.3 gives the marginal posterior distributions of the standard deviation of the Gaussian observations ($\hat{\sigma}_{\varepsilon}$), the standard deviation of the random yearly effect ($\hat{\sigma}_{\gamma}$), the range of the common spatial field $\hat{\rho}$, and the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$.
-The solid lines represent the spatial models and the dotted lines represent the non-spatial models.
-Narrower densities suggest smaller estimated uncertainty.
-We find that some of the uncertainty of the Gaussian observations $(\hat{\sigma}_{\varepsilon})$ and random yearly effect $(\hat{\sigma}_{\gamma})$ go into the spatial effect $\{\hat{\sigma}_{\omega(s)}\}$, and this can be interpreted as some of the unobserved variability (e.g., weather, temperature) being explained by the spatial rutting component; see Table 3 and Figure A.3 (a) and (b).
-This is confirmed by the smaller estimates of the ‘noise’ term $(\hat{\sigma}_{\varepsilon})$ compared with the non-spatial model (Table 4).
-We examine the effects of the range of the common spatial field ($\hat{\rho}$), and the standard deviation of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$.
-As more explanatory variables are added to model (e.g., rut depth), the variability of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$ decreases; see Table 4 and Figure A.3, graph (d).
-In Model 2, rut depth from the previous year is included, and this effect is large with the credible interval not covering zero.
-This indicates that deeper rut depths from the previous year would result in increased rutting; Figure A.2, graph (b).
-In addition, the spatial range for the common spatial field is, generally, larger for Model 2 compared with Model 3 (the simpler model), where rut depth is included (see Table LABEL:tab:range_for_each_year_appendix in Supplementary Material LABEL:supplementary:addional_results) for estimated $\hat{\rho}_{t}$ for each year).
-This is so, presumably, because of influenced factors such as varying weather patterns.
-But including lane width reduces the effect of rut depth on expected rutting.
-Refer to caption (a) Marginal posterior distributions of the standard deviation of the Gaussian observations $\hat{\sigma}_{\varepsilon}$.
-Refer to caption (b) Marginal posterior distributions of the standard deviation of the random yearly effect $\hat{\sigma}_{\gamma}$.
-Refer to caption (c) Marginal posterior distributions of the nominal range of the common spatial field $\hat{\rho}$ Refer to caption (d) Marginal posterior distributions of the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$.
+Figure A.3 presents the marginal posterior distributions of the standard deviation of Gaussian observations ($\hat{\sigma}_{\varepsilon}$), random yearly effect ($\hat{\sigma}_{\gamma}$), range of the common spatial field ($\hat{\rho}$), and standard deviation of the common spatial field ($\hat{\sigma}_{\omega(s)}$).
+Solid and dotted lines represent spatial and non-spatial models, respectively, with narrower densities indicating smaller estimated uncertainties.
+The spatial model captures some of the uncertainties in Gaussian observations and yearly effects, suggesting that unobserved variability (e.g., weather) is partly explained by the spatial rutting component.
+This is evidenced by smaller noise term estimates ($\hat{\sigma}_{\varepsilon}$) in the spatial model (Table 4).
+As more explanatory variables are added (e.g., rut depth), the variability of the spatial field ($\hat{\sigma}_{\omega(s)}$) decreases (Figure A.3, graph (d)).
+In Model 2, including previous year’s rut depth gives a large effect, indicating deeper ruts lead to increased rutting (Figure A.2, graph (b)).
+The spatial range ($\hat{\rho}$) is generally larger in Model 2 compared to Model 3, likely due to varying weather patterns, but including lane width reduces the effect of rut depth on expected rutting.
+(a) Marginal posterior distributions of the standard deviation of the Gaussian observations $\hat{\sigma}_{\varepsilon}$.
+(b) Marginal posterior distributions of the standard deviation of the random yearly effect $\hat{\sigma}_{\gamma}$.
+(c) Marginal posterior distributions of the nominal range of the common spatial field $\hat{\rho}$
+(d) Marginal posterior distributions of the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$.
 Figure A.3: Estimates of marginal posterior distributions of the spatial models (solid lines) and non-spatial models (dotted lines).
@@
 Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or go to the next url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@337039,7037105,18/vegsystemreferanse:337039.29:7037149.25
+A.5 Average mean air surface temperature
+In Norway, the temperature rarely exceeds 30 ${}^{\circ}\mathrm{C}$, and on average, monthly temperatures do not exceed 20 ${}^{\circ}\mathrm{C}$.
+This limited variation in high temperatures suggests that temperature may not play a critical role in the rutting process within this specific climatic context.
+The maximum average mean temperature per month or over years in Norway is less than 18 ${}^{\circ}\mathrm{C}$ (and less than $13.5\ ^{\circ}\mathrm{C}$ for all years, except 2104 and 2018).
+This indicates that temperature extremes, which could significantly influence rutting, are infrequent.
+Figure A.5: Average mean air surface temperature (${}^{\circ}\mathrm{C}$,  ) for years 2011-2020.
+The black circles ($\bullet$) gives the average mean monthly air surface temperature and the coloured dots indicated in the legend gives the mean monthly for each year.
+The dashed black line gives a temperature value of $13.5^{\circ}\mathrm{C}$ (see https://climateknowledgeportal.worldbank.org/country/norway/trends-variability-historical for further details).